### v2

# Counter-examples to Gamma conjecture I

Abstract.
We investigate Gamma conjecture I and its underlying Conjecture $\mathcal{O}$ for the $\mathbb{P}^{1}$-bundles $X_{n}=\mathbb{P}_{\mathbb{P}^{n}}(\mathcal{O}\oplus\mathcal{O}(n))$ with $n\geq 3$. We show that Conjecture $\mathcal{O}$ does not hold if $n$ is odd, and that Gamma conjecture I does not hold if $n$ is even. Led by this example, we propose modifications for Gamma conjecture I, discuss Gamma conjecture I over the Kähler moduli space, and identify the corresponding principal asymptotic class.
Key words and phrases:
Conjecture $\mathcal{O}$. Gamma conjecture I. Mirror symmetry. Toric Fano manifolds.
2010 Mathematics Subject Classification. Primary 14N35. Secondary 14J45, 14J33.
1. Introduction
1.1. Conjecture ${\mathcal{O}}$ and Gamma conjecture I
The quantum cohomology ring $QH(X)=(H^{*}(X)\otimes\mathbb{C}[\mathbf{q}],\star)$ of a Fano manifold $X$ is a deformation of the classical cohomology ring $H^{*}(X)=H^{*}(X,\mathbb{C})$ by incorporating genus-zero, three-pointed Gromov-Witten invariants of $X$. The quantum multiplication by the first Chern class $c_{1}(X)$ of $X$ induces a linear operator $\hat{c}_{1}=c_{1}(X)\star_{\mathbf{q}=\mathbf{1}}$ on the even part $H^{\bullet}(X):=H^{\rm ev}(X)=QH^{\rm ev}(X)|_{\mathbf{q}=\mathbf{1}}$, which is a vector space of finite dimension. In [GGI], Galkin, Golyshev and Iritani introduced Property $\mathcal{O}$ (Definition 2.1) which says
  • (1)
    the spectral radius $\rho=\max\{|u|:u\in\operatorname{Spec}(\hat{c}_{1})\}$ is itself an eigenvalue of $\hat{c}_{1}$ whose multiplicity as a root of the characteristic polynomial is one;
  • (2)
    if $u\in\operatorname{Spec}(\hat{c}_{1})$ satisfies $|u|=\rho$, then $u=e^{2\pi{\mathbf{i}}k/i_{X}}\rho$ for some $k\in{\mathbb{Z}}$,
where $\operatorname{Spec}(\hat{c}_{1})$ is the set of eigenvalues of $\hat{c}_{1}$ and $i_{X}$ is the Fano index of $X$. They also made the following conjecture with the name $\mathcal{O}$, indicating the relationship between the eigenvalue $\rho$ and the structure sheaf ${\mathcal{O}}$ via homological mirror symmetry.
Conjecture ${\mathcal{O}}$ (Conjecture 3.1.2 of [GGI]).
Every Fano manifold satisfies Property $\mathcal{O}$.
It is an important problem to study the distribution of eigenvalues of $\hat{c}_{1}$. For instance, the monotone Fukaya category of $X$ decomposes into components indexed by eigenvalues of $\hat{c}_{1}$ (see e.g. [Au, She]).
The Gamma class $\widehat{\Gamma}_{X}$ [HKTY, Libg, Iri09, Lu] of $X$ is a real characteristic class, and can be regarded as a ”square root” of the Todd class of $X$. It is defined by the Chern roots of the tangent bundle $TX$ of $X$ and Euler’s $\Gamma$-function, and has the following expansion:
$\widehat{\Gamma}_{X}=\exp\big{(}-C_{\rm eu}c_{1}(X)+\sum_{k=2}^{\infty}(-1)^{k}(k-1)!\zeta(k)\operatorname{ch}_{k}(TX)\big{)}\in H^{*}(X,\mathbb{R})$
where $C_{\rm eu}=\lim_{n\to\infty}{(-\log n+\sum_{k=1}^{n}{1\over k})}$ is the Euler-Mascheroni constant, $\zeta(k)=\sum_{n=1}^{\infty}{1\over n^{k}}$ is the value of Riemann zeta function at $k$, and $\operatorname{ch}_{k}$ denotes the $k$-th Chern character. Gamma conjectures I and II were both proposed in [GGI], which relate the quantum cohomology of $X$ and the Gamma class $\widehat{\Gamma}_{X}$ in terms of quantum differential equations. In this paper, we will focus on Gamma conjecture I, of which Property $\mathcal{O}$ is an underlying assumption.
There is a quantum connection, $\nabla_{z\partial_{z}}=z\partial_{z}-{1\over z}\hat{c}_{1}+\mu$, defined on $H^{\bullet}(X)\otimes\mathbb{C}[z,z^{-1}]$. It has a regular singularity at $z=\infty$ and an irregular singularity at $z=0$. Near $z=\infty$, there is a canonical correspondence between $H^{*}(X)$ and the space of flat sections [Giv96], and near $z=0$, the space of flat sections (along a sector) has a natural filtration by exponential growth order. It is shown in [GGI, Proposition 3.3.1] that the vector space of flat sections $s(z)$ with the smallest asymptotics as $z\to 0$ along $\mathbb{R}_{>0}$ is one-dimensional, provided that Property $\mathcal{O}$ holds. Such one-dimensional space $\mathbb{C}s(z)$ near $z=\infty$ corresponds to a class $A_{X}\in H^{*}(X)$ up to a scalar, called the principal asymptotic class of $X$.
Gamma conjecture I ([GGI, Conjecture 3.4.3]).
Let $X$ be a Fano manifold satisfying Property $\mathcal{O}$. The principal asymptotic class $A_{X}$ is given by $\widehat{\Gamma}_{X}$.
The principal asymptotic class $A_{X}$ also appears in an asymptotic expansion of Givental’s $J$-function $J_{X}(\tau,z)$ restricted to the anti-canonical line $\tau\in{\mathbb{R}}c_{1}(X)$. As explained in [GaIr, Proposition 3.8], $J_{X}(c_{1}(X)\log t,1)$ admits an asymptotic expansion at $t\to+\infty$ with leading term $A_{X}$, provided that Property $\mathcal{O}$ holds. As a consequence in [GGI, Corollary 3.6.9], Gamma conjecture I is equivalent to the limit formula:
(1) $\lim_{t\to+\infty}[J_{X}(c_{1}(X)\log t,1)]=[\widehat{\Gamma}_{X}]$
in the projective space ${\mathbb{P}}(H^{\bullet}(X))$.
Gamma conjecture I, together with its underlying Conjecture $\mathcal{O}$, has been proved for complex Grassmannians [GGI], Fano 3-fold of Picard number one [GoZa], Fano complete intersections in projective spaces [GaIr, SaSh, Ke], and del Pezzo surfaces [HKLY]. Conjecture $\mathcal{O}$ has also been proved for flag varieties $G/P$ [ChLi] by using Perron-Frobenius theorem on irreducible nonnegative matrices, as well as for a few more cases [GaGo, Che, LMS, BFSS, Wit, HKLS].
1.2. Gamma conjecture and mirror symmetry
Gamma conjectures and mirror symmetry are closely related. There are recent studies of mirror symmetry version of Gamma conjectures for Fano manifolds and Calabi-Yau manifolds [AGIS, Wang, FWZ, Iri23c]. The case for Calabi-Yau manifolds originates from [Hoso, Conjecture 2.2] by Hosono. There have been extensive studies of mirror symmetry for toric Fano manifolds in [Baty, Giv98, HoVa, ChOh, FOOO, Iri09, Abou, FLTZ] and by many others [^1] See also the surveys [Iri20a, Chan20] and references therein.. Roughly speaking, mirror symmetry asserts equivalences between geometric information on the so-called A-side and B-side. For a toric Fano manifold $X_{\Delta}$ of dimension $N$ on the A-side, the mirror object on the B-side is a Landau-Ginzburg model $(\check{X},f_{\Delta})$, where $f_{\Delta}\colon\check{X}=(\mathbb{C}^{\times})^{N}\to\mathbb{C}$ is a Laurent polynomial called the Landau-Ginzburg superpotential. Mirror symmetric Gamma conjecture for $X_{\Delta}$ focuses on the match between the Gamma-integral structure of quantum cohomology of $X_{\Delta}$ and the natural integral structure of oscillatory integrals associated with $f_{\Delta}$, and has been verified in [Iri09]. Building on this result, [GaIr, Theorem 6.3] showed that the (non-mirror-symmetric) Gamma conjecture I holds for $X_{\Delta}$, provided that $f_{\Delta}$ satisfies a B-analogue of Property $\mathcal{O}$ as defined in [GaIr, Condition 6.1].
Therefore it is natural to investigate the B-analogue of Conjecture ${\mathcal{O}}$ for the study of Gamma conjecture I for $X_{\Delta}$. To explain the condition, let us start with some facts in mirror symmetry. It was shown in [Baty, Giv98] (see also [ChLe]) that the quantum cohomology ring $QH(X_{\Delta})$ is isomorphic to the Jacobi ring $\operatorname{Jac}(f_{\Delta})$. Moreover, it was proved in [Au] (see also [OsTy]) that the eigenvalues of $\hat{c}_{1}$ (with multiplicities counted) coincide with the critical values of the superpotential $f_{\Delta}$. The B-analogue of Property $\mathcal{O}$ (Definition 2.8) says:
  • (a)
    every critical value $u$ of $f_{\Delta}$ satisfies $|u|\leq T_{\text{con}}$, where $T_{\text{con}}$ is the value of $f_{\Delta}$ at the conifold point ${\mathbf{x}}_{\text{con}}$ (a unique critical point of $f_{\Delta}|_{({\mathbb{R}}_{>0})^{N}}$ [Ga1]);
  • (b)
    ${\mathbf{x}}_{\text{con}}$ is a unique critical point of $f_{\Delta}$ contained in $f_{\Delta}^{-1}(T_{\text{con}})$.
When $X_{\Delta}$ is a blow-up of $\mathbb{P}^{n}$ along a linear subspace $\mathbb{P}^{r-1}$, the B-analogue of Property $\mathcal{O}$ was verified in [Yang] by using classical complex analysis.
1.3. Counter-examples to Conjecture ${\mathcal{O}}$ and Gamma conjecture I
In this paper, we investigate certain toric Fano manifolds of Picard number two. This paper’s central contribution is the following result, which summarizes Theorems 3.9, 4.5 and 4.7.
Theorem 1.1.
Let $X_{n}=\mathbb{P}_{\mathbb{P}^{n}}(\mathcal{O}\oplus\mathcal{O}(n))$, where $n\geq 3$.
  1. (1)
    If $n$ is even, then Conjecture $\mathcal{O}$ holds for $X_{n}$, while Gamma conjecture I does not hold for $X_{n}$.
  2. (2)
    If $n$ is odd, then Conjecture $\mathcal{O}$ does not hold for $X_{n}$, while a variant of Gamma conjecture I holds for $X_{n}$.
We note that both Conjecture $\mathcal{O}$ and Gamma conjecture I hold for $n\in\{1,2\}$. When $n=1$, $X$ is a blow-up of $\mathbb{P}^{2}$ at one point, and both conjectures have been verified in [HKLY]. The case for $n=2$ will be verified in Sections 3 and 4. For Conjecture $\mathcal{O}$, we use mirror symmetry and reduce the analysis of the critical values of the mirror superpotential to optimization problems in nonlinear programming. Such reduction seems effective: it also works for $\mathbb{P}_{\mathbb{P}^{n}}(\mathcal{O}\oplus\mathcal{O}(n-1))$ and a few more examples as we discuss in the appendix. This method is also applied in [HKLSu] to find a counterexample of the lower bound conjecture in [Ga1].
To explain the variant of Gamma conjecture I in the theorem, we introduce the following terminology. We say that a linear operator $A$ on a finite-dimensional vector space has a simple rightmost eigenvalue $u\in{\mathbb{C}}$ if (1) $u$ is an eigenvalue of $A$ whose multiplicity as a root of the characteristic polynomial is one and if (2) any other eigenvalue $u^{\prime}$ of $A$ satisfies $\operatorname{Re}(u^{\prime})<\operatorname{Re}(u)$. We consider the following condition (Definition 2.2):
(*) the operator $\hat{c}_{1}$ has a simple rightmost eigenvalue $\rho^{\prime}$.
Since $\hat{c}_{1}$ is a real operator, the simple rightmost eigenvalue $\rho^{\prime}$ (if it exists) is a real number and is given by $\max\{\operatorname{Re}(u):u\in\operatorname{Spec}(\hat{c}_{1})\}$. As explained in [GGI, Remark 3.1.9], when condition (*1.3) holds, the space of $\nabla$-flat sections with the smallest asymptotics as $z\to+0$ is one-dimensional, and the principal asymptotic class can still be defined. In the odd $n$ case of our theorem, condition (*1.3) holds, and the variant of Gamma conjecture I says that the principal asymptotic class is the Gamma class.
Remark 1.2.
Hugtenburg [Hug24] recently studied Gamma conjecture I for certain non-Kähler monotone symplectic manifolds constructed by Fine-Panov [FiPa]. In those examples, the spectral radius $\rho$ of $\hat{c}_{1}$ is a non-simple eigenvalue of $\hat{c}_{1}$ and Conjecture ${\mathcal{O}}$ fails. However, Gamma conjecture I still holds in the sense that the limit formula (1) is true.
The counterexample $X_{n}$ to Conjecture $\mathcal{O}$/Gamma conjecture I leads to many interesting questions. For instance for Conjecture $\mathcal{O}$, we have numerically analyzed the eigenvalues of $\hat{c}_{1}$ for all the 8635 toric Fano manifolds of dimension less than 7. As from Table 1 in Section 5.1, most of these manifolds satisfy part (1) of Property $\mathcal{O}$ (see Definition 2.1). Among the remaining cases, almost all of the manifolds satisfy the condition (*1.3). In total fewer than $4\%$ of the manifolds do not satisfy the condition (*1.3). It would be interesting to explore a geometric/combinatorial characterization of the counterexamples for general toric Fano manifolds. Regarding Gamma conjecture I, there are at least two possible directions for investigation. Firstly, we may seek to correct the original sufficient condition (which assumes Property $\mathcal{O}$ solely), so that the principal asymptotic class $A_{X}$ remains to be given by the Gamma class $\widehat{\Gamma}_{X}$. Alternatively, we may explore what $A_{X}$ is, if it is not given by $\widehat{\Gamma}_{X}$.
1.4. Exploring Gamma conjecture I over the Kähler moduli space
In the latter half of the paper, we explore Gamma conjecture I for a general quantum parameter ${\mathbf{q}}$ and determine the principal asymptotic class $A_{X}$ of the counter-example. The paramter ${\mathbf{q}}$ of quantum cohomology lies in the so-called “Kähler moduli space”
$\mathcal{M}_{X}:=\operatorname{Hom}(H_{2}(X,{\mathbb{Z}})/{\text{torsion}},{\mathbb{C}}^{\times})\cong H^{2}(X,{\mathbb{Z}})\otimes{\mathbb{C}}^{\times}.$
The original Gamma conjecture I focused on the quantum multiplication of $c_{1}(X)$ at ${\mathbf{q}}=\mathbf{1}$, but this specialization lacked a strong theoretical justification. Exploring Gamma conjecture I for general parameters ${\mathbf{q}}\in\mathcal{M}_{X}$ offers two advantages: (1) we can connect the behaviour of eigenvalues of the quantum multiplication by $c_{1}(X)$ with the birational geometry of $X$, and (2) we can sometimes determine the principal asymptotic class through mutation from the asymptotic class near a boundary of the Kähler moduli space, which is typically easier to determine.
First of all, we generalize the condition (*1.3) for the operator $\hat{c}_{1}=(c_{1}(X)\star_{{\mathbf{q}}=\mathbf{1}})$ on $H^{\bullet}(X)$ to the condition that $(c_{1}(X)\star_{\mathbf{q}})$ has a simple rightmost eigenvalue (see Definition 6.6). Assuming the existence of a simple rightmost eigenvalue $u$ for $(c_{1}(X)\star_{\mathbf{q}})$, we can define the principal asymptotic class $A_{X}(\tau)$ at $\tau\in H^{2}(X,{\mathbb{C}})$ (Proposition/Definition 6.8), where $\tau$ represents a point on the universal covering $H^{2}(X,{\mathbb{C}})\to\mathcal{M}_{X}$ lifting ${\mathbf{q}}\in\mathcal{M}_{X}$ (which we write as ${\mathbf{q}}=e^{\tau}$). The aforementioned counter-example $X_{n}$ motivates many questions, including the following in Question 6.12:
  • Does there exist a point ${\mathbf{q}}\in\mathcal{M}_{X}$ such that $(c_{1}(X)\star_{\mathbf{q}})$ has a simple rightmost eigenvalue and that $[A_{X}(\tau)]=[\widehat{\Gamma}_{X}]$?
  • Can we find such a $\tau$ within $H^{2}(X,{\mathbb{R}})$?
  • Can we characterize the region in $H^{2}(X)$ where $[A_{X}(\tau)]=[\widehat{\Gamma}_{X}]$ holds?
If $[A_{X}(\tau)]\neq[\widehat{\Gamma}_{X}]$, we can further ask about the possible forms of $A_{X}(\tau)$. Dubrovin’s conjecture [Du1] and Gamma conjecture II [GGI] suggest that there should be an exceptional object $V$ in the derived category $D^{b}(X)$ of coherent sheaves on $X$ such that
$[A_{X}(\tau)]=[\widehat{\Gamma}_{X}\operatorname{Ch}(V)]$
where $\operatorname{Ch}(V)=\sum_{p\geq 0}(2\pi{\mathbf{i}})^{p}\operatorname{ch}_{p}(V)$ is the modified Chern character, see Question 6.13. The result [Iri09, §4.3] on toric mirror symmetry implies the following result:
Theorem 1.3 (Theorem 7.7).
Let $X$ be a toric Fano manifold. Suppose that $(c_{1}(X)\star_{\mathbf{q}})$ has a simple rightmost eigenvalue $u$ for some ${\mathbf{q}}=e^{\tau}\in\mathcal{M}_{X}$. Then the principal asymptotic class $A_{X}(\tau)$ is of the form $\widehat{\Gamma}_{X}\operatorname{Ch}(V)$ for some $V$ in the topological $K$-group $K^{0}(X)$.
The spectrum of the quantum multiplication $(c_{1}(X)\star_{\mathbf{q}})$ reflects the birational geometry of $X$. When ${\mathbf{q}}$ approaches a boundary of $\mathcal{M}_{X}$ given by an extremal ray (in the sense of Mori theory), the spectrum of the quantum cohomology of $X$ clusters in a specific way predicted by the corresponding extremal contraction. It is therefore easier to study Gamma conjecture I when ${\mathbf{q}}$ is close to such a boundary.
We study this connection to birational geometry in detail for the example $X_{n}={\mathbb{P}}_{{\mathbb{P}}^{n}}({\mathcal{O}}\oplus{\mathcal{O}}(n))$. The space $X_{n}$ has two extremal contractions: the divisorial contraction $\varphi\colon X_{n}\to{\mathbb{P}}(1,\dots,1,n)$ to the weighted projective space (where $1$ appears $n+1$ times) and the ${\mathbb{P}}^{1}$-fibration $\pi\colon X_{n}\to{\mathbb{P}}^{n}$. Each contraction corresponds to an extremal ray of the Mori cone and associates a line in the boundary of $\mathcal{M}_{X}$. The distributions of eigenvalues of $(c_{1}(X)\star_{\mathbf{q}})$ near these boundaries are shown in Figures 5 and 6. The principal asymptotic classes near these boundaries can be described as follows:
Theorem 1.4 (Theorems 7.8, 7.10).
Let $n\geq 2$ be even. Let ${\mathbf{q}}$ be a positive real quantum parameter and let $\tau\in H^{2}(X_{n},{\mathbb{R}})$ be the corresponding real class. Then the principal asymptotic class $A_{X_{n}}(\tau)$ near each boundary is given respectively by
$\begin{cases}\widehat{\Gamma}_{X_{n}}&\text{for ${\mathbf{q}}$ near the ``${\mathbb{P}}^{1}$-fibration boundary''}\\ \widehat{\Gamma}_{X_{n}}\operatorname{Ch}({\mathcal{O}}_{E}(-n/2))&\text{for ${\mathbf{q}}$ near the ``divisorial contraction boundary''}\end{cases}$
where $E\cong{\mathbb{P}}^{n}$ is the exceptional divisor of $\varphi$. Furthermore, if $n$ is even and $n\geq 4$, we have $A_{X_{n}}(0)=\widehat{\Gamma}_{X_{n}}\operatorname{Ch}({\mathcal{O}}_{E}(-n/2))$.
In this theorem, we derive the principal asymptotic class near the “divisorial contraction boundary” from the one near the “${\mathbb{P}}^{1}$-fibration boundary” through a sequence of mutation.
History and organization of the paper
In an early stage of this project, J. Hu, H. Ke, C. Li and Z. Su attempted to prove Conjecture $\mathcal{O}$ and Gamma conjecture I for toric Fano manifolds of Picard number two. However, they discovered counterexamples and authored Sections 3-5 of the current paper. In later communications, S. Galkin and H. Iritani suggested considering deformations over the Kähler moduli space and identified the principal asymptotic class corresponding to the unexpected eigenvalue. They contributed Sections 6-7 of the current paper.
This paper is organized as follows. In Section 2, we describe the precise statements of conjecture $\mathcal{O}$ and Gamma conjecture I, and review basic facts of mirror symmetry for Fano manifolds. In Section 3, we study Conjecture $\mathcal{O}$ for $\mathbb{P}_{\mathbb{P}^{n}}(\mathcal{O}\oplus\mathcal{O}(n))$, by a reduction to optimization problems in nonlinear programming. In Section 4, we investigate Gamma conjecture I for $\mathbb{P}_{\mathbb{P}^{n}}(\mathcal{O}\oplus\mathcal{O}(n))$. In Section 5, we discuss possible modifications of Gamma conjecture I. In Sections 6 and 7, we discuss the Gamma conjecture I over the Kähler moduli space, demonstrating a close relationship to the birational geometry of the Fano manifold. In Section 6, we introduce the principal asymptotic class $A_{X}(\tau)$ of a Fano manifold $X$ for a general $\tau\in H^{2}(X,{\mathbb{C}})$; in Section 7 we compute it for the example ${\mathbb{P}}_{{\mathbb{P}}^{n}}({\mathcal{O}}\oplus{\mathcal{O}}(n))$ on certain regions of the Kähler moduli space. Finally in the appendix, we study Conjecture $\mathcal{O}$ for $\mathbb{P}_{\mathbb{P}^{n}}(\mathcal{O}\oplus\mathcal{O}(n-1))$.
Acknowledgements
The authors would like to thank Kwok Wai Chan, Bohan Fang, Kai Hugtenburg, King-Leung Lee, Jiayu Song and Heng Xie for helpful discussions, and to thank Zhi Chen for collecting the toric data from [Obro2] for toric Fano manifolds of dimensions less than 7. Z. Su would like to thank Xiaowei Wang for constant encouragement. J. Hu, C. Li and H. Ke are supported in part by the National Key Research and Development Program of China No. 2023YFA100980001. H. Ke is also supported in part by NSFC Grant 12271532. S. Galkin is supported by CNPq grant PQ 315747, and Coordenação de Aperfeiçoamento de Pessoal de Nível Superior-Brasil (CAPES)-Finance Code 001. H. Iritani is supported by JSPS grant 21H04994 and 23H01073.
2. Preliminaries
2.1. (Variant of) Gamma conjecture I for Fano manifolds
In this section, we briefly review Conjecture ${\mathcal{O}}$ and Gamma conjecture I for Fano manifolds following [GGI, GaIr] and [HKLY, Section 2.1]. We also explain a variant of Gamma conjecture I that only assumes the weaker condition (*1.3).
2.1.1. Quantum cohomology
We refer to [CoKa] for background material on Gromov-Witten invariants and quantum cohomology.
Let $X$ be a Fano manifold, namely a compact complex manifold $X$ whose anticanonical line bundle is ample. Let ¯$\overline{\mathcal{M}}_{0,k}(X,\mathbf{d})$ denote the moduli stack of $k$-pointed genus-zero stable maps $(f:C\to X;p_{1},...,p_{k})$ of class $\mathbf{d}\in H_{2}(X,\mathbb{Z})$, which has a coarse moduli space $\overline{M}_{0,k}(X,\mathbf{d})$. Let $[\overline{M}_{0,k}(X,\mathbf{d})]^{\rm virt}$ be the virtual fundamental class of ¯$\overline{\mathcal{M}}_{0,k}(X,\mathbf{d})$, which is of complex degree $\dim X-3+\int_{\mathbf{d}}c_{1}(X)+k$ in the Chow group $A_{*}(\overline{\mathcal{M}}_{0,k}(X,\mathbf{d}))$. Given classes $\gamma_{1},...,\gamma_{k}\in H^{*}(X)=H^{*}(X,\mathbb{C})$ and nonnegative integers $a_{i}$ for $1\leq i\leq k$, we have the following associated gravitational correlator
$\langle\psi^{a_{1}}\gamma_{1},...,\psi^{a_{k}}\gamma_{k}\rangle_{0,k,\mathbf{d}}:=\int_{[\overline{M}_{0,k}(X,\mathbf{d})]^{\rm virt}}\prod_{i=1}^{k}\big{(}c_{1}(\mathcal{L}_{i})^{a_{i}}\cup\operatorname{ev}_{i}^{*}(\gamma_{i})\big{)}.$
Here $\mathcal{L}_{i}$ denotes the line bundle on ¯$\overline{\mathcal{M}}_{0,k}(X,\mathbf{d})$ whose fiber over the stable map $(f:C\to X;p_{1},...,p_{k})$ is the cotangent space $T^{*}_{p_{i}}C$, and $\operatorname{ev}_{i}$ denotes the $i$-th evaluation map.
The (small) quantum product of $\alpha,\beta\in H^{*}(X)$ is defined by
(2) $\alpha\star_{\mathbf{q}}\beta:=\sum_{\mathbf{d}\in H_{2}(X,\mathbb{Z})}\sum_{i}\langle\alpha,\beta,\phi_{i}\rangle_{0,3,\mathbf{d}}\phi^{i}\mathbf{q}^{\mathbf{d}}.$
Here $\{\phi_{i}\}$ is a homogeneous basis of $H^{*}(X)$, $\{\phi^{i}\}$ is its dual basis in $H^{*}(X)$ with respect to the Poincaré pairing; the quantum parameter $\mathbf{q}$ lies in the “Kähler moduli space”
$\mathcal{M}_{X}:=\mathrm{Hom}(H_{2}(X,{\mathbb{Z}})/\mathrm{torsion},{\mathbb{C}}^{\times}),$
and the monomial function $\mathbf{q}^{\mathbf{d}}\colon\mathcal{M}_{X}\to{\mathbb{C}}$ is the evaluation at $\mathbf{d}\in H_{2}(X,{\mathbb{Z}})$. Note that $\mathbf{q}^{\mathbf{d}}$ has a non-zero coefficient in (2) only if $\mathbf{d}$ is effective. By the Fano condition, there are only finitely many $\mathbf{d}$ such that $\mathbf{q}^{\mathbf{d}}$ has a non-zero coefficient. The quantum product $\star_{\mathbf{q}}$ is a family of products on $H^{*}(X)$ with parameter $\mathbf{q}$, and it is a deformation of the cup product:
$\lim\limits_{\mathbf{q}\to 0}\alpha\star_{\mathbf{q}}\beta=\alpha\cup\beta.$
2.1.2. Conjecture ${\mathcal{O}}$ and condition (*1.3)
Consider the even part of the cohomology $H^{\bullet}(X):=H^{\rm even}(X)$ and the finite-dimensional $\mathbb{C}$-algebra $QH^{\bullet}(X)=(H^{\bullet}(X),\bullet)$ with the product defined by $\alpha\bullet\beta:=\alpha\star_{\mathbf{q}=\mathbf{1}}\beta$, namely by specialization of the quantum parameter to the constant function with value $1$. Let $\hat{c}_{1}$ denote the linear operator induced by the first Chern class:
$\hat{c}_{1}:H^{\bullet}(X)\longrightarrow H^{\bullet}(X);\,\beta\mapsto c_{1}(X)\bullet\beta.$
Definition 2.1 (Property ${\mathcal{O}}$).
For a Fano manifold $X$, we denote by $\rho=\rho(\hat{c}_{1})$ the spectral radius of the linear operator $\hat{c}_{1}$, namely
$\rho:=\max\{|\lambda|:\lambda\in\operatorname{Spec}(\hat{c}_{1})\}$
where $\operatorname{Spec}(\hat{c}_{1}):=\{\lambda:\lambda\in{\mathbb{C}}\mbox{ is aneigenvalue of }\hat{c}_{1}\}$. We say that $X$ satisfies Property ${\mathcal{O}}$ if the following two conditions are satisfied.
  1. (1)
    $\rho$ is a simple root of the characteristic polynomial of $\hat{c}_{1}$ (in particular $\rho\in\operatorname{Spec}(\hat{c}_{1})$);
  2. (2)
    for any $\lambda\in\operatorname{Spec}(\hat{c}_{1})$ with $|\lambda|=\rho$, we have $\lambda^{i_{X}}=\rho^{i_{X}}$, where $i_{X}$ denotes the Fano index of $X$ defined by $i_{X}:=\max\{k\in{\mathbb{Z}}~{}:~{}{c_{1}(X)\over k}\in H^{2}(X,\mathbb{Z})\}$.
Conjecture ${\mathcal{O}}$ (Conjecture 3.1.2 of [GGI]).
Every Fano manifold satisfies Property ${\mathcal{O}}$.
Conjecture ${\mathcal{O}}$ has proven true for any flag variety $G/P$ in [ChLi] by using Perron-Frobenius theorem based on a remark due to Kaoru Ono.
We also recall condition (*1.3) from the Introduction, which is a weaker version of Property ${\mathcal{O}}$. As remarked in [GGI, Remark 3.1.9], we can study Gamma conjecture I under this weaker condition.
Definition 2.2 (Condition (*)).
We say that $\hat{c}_{1}$ has a simple rightmost eigenvalue $u\in{\mathbb{C}}$ if
  • (1)
    $u$ is a simple root of the characteristic polynomial of $\hat{c}_{1}$ and
  • (2)
    any other eigenvalues $u^{\prime}$ of $\hat{c}_{1}$ satisfies $\operatorname{Re}(u^{\prime})<\operatorname{Re}(u)$.
We say that $X$ satisfies condition (*1.3) if $\hat{c}_{1}$ has a simple rightmost eigenvalue. In this case, the simple rightmost eigenvalue is given by
$\rho^{\prime}:=\max\{\operatorname{Re}(u):u\in\operatorname{Spec}(\hat{c}_{1})\}$
as $\hat{c}_{1}$ is a real operator. Note that $\rho^{\prime}=\rho$ when Property ${\mathcal{O}}$ holds.
Remark 2.3.
Here we regard $\hat{c}_{1}$ as an operator on the even part $H^{\bullet}(X)$. By the arguments in [SaSh, GaIr], we find that Property ${\mathcal{O}}$ or condition (*1.3) remains equivalent even if we consider $\hat{c}_{1}$ as an opeartor on the entire cohomology group $H^{*}(X)$ or on the smaller part $\bigoplus_{p}H^{p,p}(X)\subset H^{\bullet}(X)$.
2.1.3. (Variant of) Gamma conjecture I
There is a quantum connection on the trivial $H^{\bullet}(X)$-bundle over $\mathbb{P}^{1}$, given by
$\nabla_{z\partial_{z}}=z\partial_{z}-{1\over z}(c_{1}(X)\bullet)+\mu.$
Here $z$ is an inhomogeneous co-ordinate on $\mathbb{P}^{1}$ and $\mu\in\operatorname{End}(H^{\bullet}(X))$ is the Hodge grading operator defined by $\mu(\phi)=(p-{\dim X\over 2})\phi$ for $\phi\in H^{2p}(X)$. The quantum connection is a meromorphic connection, which is logarithmic at $z=\infty$ and irregular singular at $z=0$. The space of flat sections can be identified with the cohomology group $H^{\bullet}(X)$ via the fundamental solution
$S(z)z^{-\mu}z^{c_{1}}$
where $S\colon{\mathbb{P}}^{1}\setminus\{0\}\to\operatorname{End}(H^{\bullet}(X))$ is a holomorphic function given by
$S(z)\alpha=\alpha+\sum_{i}\sum_{{\mathbf{d}}\neq 0}\left\langle\frac{\alpha}{-z-\psi},\phi^{i}\right\rangle_{0,2,{\mathbf{d}}}\phi_{i}$
and $c_{1}\in\operatorname{End}(H^{\bullet}(X))$ is the cup product operator by $c_{1}(X)$. Here, $\alpha/(-z-\psi)$ should be expanded in the series $\sum_{n\geq 0}\alpha\psi^{n}(-z)^{-n-1}$. See [GGI, Proposition 2.3.1] for detailed explanations.
Proposition/Definition 2.4 ([GGI, Proposition 3.3.1, Remark 3.1.9]).
If $X$ satisfies Property ${\mathcal{O}}$, or even more weakly, if $X$ satisfies condition (*1.3), then the space
$\{\alpha\in H^{\bullet}(X):e^{\rho^{\prime}/z}S(z)z^{-\mu}z^{c_{1}}(\alpha)\mbox{ has moderate growth as }z\to+0\}$
is a one-dimensional linear subspace of $H^{\bullet}(X)$, where ‘moderate growth’ means that $\|e^{\rho^{\prime}/z}S(z)z^{-\mu}z^{c_{1}}(\alpha)\|=O(z^{-m})$ as $z\to+0$ for some $m\in\mathbb{N}$. The principal asymptotic class $A_{X}$ is a generator of this space defined up to scalar.
[…]
$\widehat{\Gamma}_{X}:=\prod_{i=1}^{n}\Gamma(1+\delta_{i})=\exp\big{(}-C_{\rm eu}c_{1}(X)+\sum_{k=2}^{\infty}(-1)^{k}(k-1)!\zeta(k)\operatorname{ch}_{k}(TX)\big{)}\in H^{\bullet}(X,\mathbb{R})$
where $C_{\rm eu}$ is the Euler-Mascheroni constant, $\zeta(k)=\sum_{n=1}^{\infty}{1\over n^{k}}$ is the value of Riemann zeta function at $k$, and $\operatorname{ch}_{k}$ denotes the $k$-th Chern character.
Gamma conjecture I ([GGI, Conjecture 3.4.3]).
Let $X$ be a Fano manifold satisfying Property ${\mathcal{O}}$. Then the principal asymptotic class $A_{X}$ is given by the Gamma class $\widehat{\Gamma}_{X}$.
Variant of Gamma conjecture I.
Let $X$ be a Fano manifold satisfying condition (*1.3). Then the principal asymptotic class $A_{X}$ is given by the Gamma class $\widehat{\Gamma}_{X}$.
The principal asymptotic class also appears in an asymptotic expansion of Givental’s $J$-function. Givental’s $J$-function is a cohomology-valued function in $(\tau,z)\in H^{\bullet}(X)\times{\mathbb{C}}^{\times}$ given by descendant Gromov-Witten invariants
(3) $J_{X}(\tau,z)=1+\frac{\tau}{z}+\sum_{i}\sum_{(n,{\mathbf{d}})\neq(0,0)}\left\langle 1,\tau,\cdots,\tau,\frac{\phi^{i}}{z-\psi}\right\rangle_{0,n+2,{\mathbf{d}}}\phi_{i}$
where $\phi^{i}/(z-\psi)$ should be expanded in the series $\sum_{n\geq 0}\phi^{i}\psi^{n}z^{-n-1}$. The restriction to the anticanonical line $\tau=c_{1}(X)\log t$ and $z=1$ is given in terms of the fundamental solution $S(z)z^{-\mu}z^{c_{1}(X)}$ as follows:
$J_{X}(c_{1}(X)\log t,1)=z^{\frac{\dim X}{2}}\big{(}S(z)z^{-\mu}z^{c_{1}}\big{)}^{-1}1$
with $t=z^{-1}$ (see also [HKLY, Remark 2.2]).
The following result follows from the same argument in [GaIr, Proposition 3.8], which is based on [GGI, Propositions 3.6.2, 3.2.1].
Proposition 2.5.
For any Fano manifold $X$ satisfying condition (*1.3), Givental’s $J$-function $J_{X}(c_{1}(X)\log t,1)$ has an asymptotic expansion of the form
(4) $J_{X}(c_{1}(X)\log t,1)\sim Ct^{-{\dim X\over 2}}e^{\rho^{\prime}t}(A_{X}+\alpha_{1}t^{-1}+\alpha_{2}t^{-2}+\cdots)$
as $t\to+\infty$ on the positive real line, where $C$ is a non-zero constant and $\alpha_{i}\in H^{\bullet}(X)$.
As a consequence, an equivalent description of (variant of) Gamma conjecture I is given in terms of the $J$-function $J_{X}(c_{1}(X)\log t,1)$ as follows (cf. [GGI, Corollray 3.6.9]):
Corollary 2.6.
Let $X$ be a Fano manifold satisfying Property ${\mathcal{O}}$ (or condition (*1.3)). Then Gamma conjecture I (resp. variant of Gamma conjecture I) holds for $X$ if and only if
$[\widehat{\Gamma}_{X}]=\lim_{t\to\infty}[J_{X}(c_{1}(X)\log t,1)].$
holds in the projective space ${\mathbb{P}}(H^{\bullet}(X))$ of the cohomology group.
2.2. Mirror symmetry for toric Fano manifolds
We refer readers to [CLS] for detailed account of toric varieties. A toric variety $X$ of complex dimension $N$ is an algebraic variety containing an algebraic torus as an open dense subset, such that the action of the torus on itself extends to the whole variety. Geometric information of $X=X_{\Sigma}$ can be determined by the combinatorics of its associated fan $\Sigma$ in ${\mathbb{R}}^{N}$. Here we always assume $X_{\Sigma}$ to be smooth and Fano, namely $\Sigma=\Sigma(\Delta)$ is the normal fan of a convex polytope $\Delta$ with nice properties. Thus $X_{\Sigma}$ is also dentoed as $X_{\Delta}$ alternatively.
Let $f=f_{\Sigma}=f_{\Delta}$ denote the Landau-Ginzburg superpotential mirror to the toric Fano manifold $X_{\Sigma}=X_{\Delta}$. It is a Laurent polynomial, which can be read off immediately from $\Sigma$. Precisely, we denote by $b_{1},...,b_{m}\in{\mathbb{Z}}^{N}$ the primitive generators of rays in $\Sigma$. Then the Picard number of $X_{\Sigma}$ is equal to $m-N$, and $f$ is given by
(5) $f:({\mathbb{C}}^{\times})^{N}\rightarrow{\mathbb{C}};\,\mathbf{x}\mapsto f(\mathbf{x})=\mathbf{x}^{b_{1}}+\cdots+\mathbf{x}^{b_{m}},$
where $\mathbf{x}:=(x_{1},...,x_{N})$ and for $b_{i}=(b_{i1},...,b_{iN})$, $\mathbf{x}^{b_{i}}:=x_{1}^{b_{i1}}\cdots x_{N}^{b_{iN}}$.
Mirror symmetry between quantum cohomology of $X_{\Sigma}$ and oscillatory integrals of $f$ has been proved in [Giv95, Giv98, Iri09]. As a remarkable property, there exists an isomorphism $\Psi$ of $\mathbb{C}$-algebras between quantum cohomology ring and Jacobian ring,
$\Psi\colon QH(X)|_{\mathbf{q}=1}\xrightarrow{\cong}\operatorname{Jac}(f)={\mathbb{C}}[x_{1}^{\pm 1},...,x_{N}^{\pm 1}]/(x_{1}\partial_{x_{1}}f,...,x_{N}\partial_{x_{N}}f).$
Here we set $\mathbf{q}=\mathbf{1}$ to introduce $f$ without introducing any deformations. Moreover, we have
Proposition 2.7 ([OsTy, Corollary G]).
We have $\Psi(c_{1}(X))=[f]$. Specifically, the eigenvalues of $\hat{c}_{1}$, including their multiplicities, coincide with the critical values of $f$.
Here we notice that eigenvalues of the linear operator on $\operatorname{Jac}(f)$ induced by the function multiplication by $f$ are precisely the critical values of $f$.
We can use oscillatory integrals of $f$ to get information of quantum cohomology of $X_{\Sigma}$. For example, it follows from [Iri09, Theorem 4.14] that
(6) $\langle{1},S(z)z^{-\mu}z^{c_{1}}\widehat{\Gamma}_{X}\rangle^{X}=z^{-\frac{N}{2}}\int_{({\mathbb{R}}_{>0})^{N}}e^{-\frac{f(\mathbf{x})}{z}}\frac{dx_{1}\cdots dx_{N}}{x_{1}\cdots x_{N}}.$
The restriction $f|_{({\mathbb{R}}_{>0})^{N}}$ is a real function on $({\mathbb{R}}_{>0})^{N}$ that admits a global minimum at a unique point $\mathbf{x}_{\rm con}\in({\mathbb{R}}_{>0})^{N}$ [Ga1, GGI, GaIr]. Such point $\mathbf{x}_{\rm con}$ is called the conifold point of $f$. Let $T_{\rm con}:=f(\mathbf{x}_{\rm con})$.
Definition 2.8 ([GaIr, Condition 6.1]).
We say that $X_{\Sigma}$ satisfies the B-analogue of Property ${\mathcal{O}}$, if its mirror superpotential $f=f_{\Sigma}$ satisfies the following two conditions:
  1. (a)
    every critical value $u$ of $f$ satisfies $|u|\leq T_{\rm con}$;
  2. (b)
    the conifold point $\mathbf{x}_{\rm con}$ is the unique critical point of $f$ contained in $f^{-1}(T_{\rm con})$.
Proposition 2.9 ([GaIr, Theorem 6.3]).
Suppose that a toric Fano manifold $X_{\Sigma}$ satisfies the B-analogue of Property ${\mathcal{O}}$. Then $X_{\Sigma}$ satisfies Gamma conjecture I.
3. Conjecture $\mathcal{O}$ for $X_{n}$
In the rest of this paper, we will mainly investigate the example
$X_{n}:=\mathbb{P}_{\mathbb{P}^{n}}(\mathcal{O}\oplus\mathcal{O}(n)).$
Noting $\mathbb{P}_{\mathbb{P}^{n}}(\mathcal{O}\oplus\mathcal{O}(n))\cong\mathbb{P}_{\mathbb{P}^{n}}(\mathcal{O}\oplus\mathcal{O}(-n))$, we always assume $n\geq 1$. In this section, we analyze the eigenvalues of $\hat{c}_{1}$ for $X_{n}$ via mirror symmetry. As we will show in Theorem 3.9, Conjecture $\mathcal{O}$ does not hold for $X_{n}$ when $n>1$ is odd.
3.1. Notations
The variety $X_{n}$ is a toric Fano manifold of Picard number two and of Fano index one. Let $e_{i}:=(0,\ldots,0,1,0,\ldots,0)$ (with $1$ in the $i$th position), $11\leq i\leq n+$ 1 denote the standard basis vectors of $\mathbb{R}^{n+1}$. There are exactly $(n+3)$ primitive ray generators $b_{i}$ of the associated fan of the toric variety $X_{n}$, given by
(7) $b_{i}$ $=e_{i},\quad\mbox{for }1\leq i\leq n+1;\quad b_{n+2}=-\sum_{i=1}^{n}e_{i}+ne_{n+1},\quad b_{n+3}=-e_{n+1}.$
Therefore the mirror superpotential $f$ defined on (×$(\mathbb{C}^{\times})^{n+1}$ in (5) becomes
(8) $f(\textbf{x})=x_{1}+x_{2}+\cdots+x_{n}+x_{n+1}+\frac{x_{n+1}^{n}}{x_{1}x_{2}\cdots x_{n}}+\frac{1}{x_{n+1}}.$
We adopt the following notation to analyze the critical values of $f$.
$D_{R}$ :=$:=\{t\in\mathbb{C}:|t|\leq R\},\quad\partial D_{R}:=\{t\in\mathbb{C}:|t|=R\},\quad\mathring{D}_{R}:=D_{R}\setminus\partial D_{R};$
$N_{[a,b]}$ :=$:=\{t\in\mathbb{C}:a\leq|t|\leq b\},\quad\mbox{where }b>a>0.$
Namely $D_{R}$ is the closed disc in $\mathbb{C}$ of radius $R>0$ centred at the origin, with $\partial D_{R}$ being its boundary; $N_{[a,b]}$ is an annulus in $\mathbb{C}$.
3.2. Critical values of the superpotential
We assume $n\geq 2$, since $X_{1}$ is the blow-up of $\mathbb{P}^{2}$ at a point, and has been studied in [HKLY].
[…]
$x_{1}=x_{2}=\cdots=x_{n}=:x\in\mathbb{C}^{\times}.$
Denote $y:=x_{n+1}\in{\mathbb{C}}^{\times}$. Then the system of equations $\frac{\partial f}{\partial x_{i}}=0$ can be reduced to
(9) $\begin{cases}1-\dfrac{y^{n}}{x^{n+1}}=0,\\ 1+n\dfrac{y^{n-1}}{x^{n}}-\dfrac{1}{y^{2}}=0,\end{cases}\quad(x,y)\in({\mathbb{C}}^{\times})^{2}.$
Clearly, each $t\in\mathbb{C}^{\times}$ gives a solution $(t^{n},t^{n+1})$ to the first equation in (9), and conversely, every solution $(x,y)$ to this equation admits such parametrization by simply taking $t={y\over x}$. Therefore the system (9) is equivalent to
(10) $t^{2n+2}+nt^{2n+1}-1=0.$
Consequently, the critical values of $f$ (with multiplicity counted) are precisely given by
$g(t):=nt^{n}+t^{n+1}+t^{n}+\frac{1}{t^{n+1}}$
with respect to the $22n+$ 2 roots of (10) .
Now let us move to the following optimization problem in nonlinear programming.
Problem 3.1.
$\begin{array}[]{ll@{}ll}\textnormal{maximize}&\big{|}&t^{n+1}+(n+1)t^{n}+\frac{1}{t^{n+1}}\big{|}&\\ \\ \textnormal{subject to}&&t^{2n+2}+nt^{2n+1}=1,\quad t\in\mathbb{C}^{\times}.\end{array}$
The above problem aims to find one $t_{0}$ such that $|g(t_{0})|=\max|g(t)|$ among those $t$ with the constraint. Clearly, this is quite closely related with Property $\mathcal{O}$. Indeed, $X_{\Delta}$ satisfies part (1) of Property $\mathcal{O}$ if and only if the aforementioned $t_{0}$ is unique and satisfies $g(t_{0})\in\mathbb{R}_{\geq 0}$. Denote
(11) $h(t):=t^{2n+2}+nt^{2n+1},\quad\tilde{g}(t):=t^{n}+\frac{2}{t^{n+1}},\quad\check{g}(t):=2t^{n+1}+(2n+1)t^{n}.$
Notice $h(t)=1$ if and only if $t^{n+1}+nt^{n}=\frac{1}{t^{n+1}}$. Whenever this holds, we have
$g(t)=\tilde{g}(t)=\check{g}(t).$
To maximize $|g(t)|$, we analyze the distribution of the roots of $h(t)-1$ as follows.
One can check that $h(t)-1$ and $h^{\prime}(t)$ have no common roots, implying that they are coprime. So every root of $h(t)-1$ has multiplicity one. Viewed as a real function, $h(t)$ is decreasing on $(-\infty,-\frac{n(2n+1)}{2n+2})$ and increasing on $(-\frac{n(2n+1)}{2n+2},\infty)$, possessing exactly two critical points at $t=0$ and $t=-\frac{n(2n+1)}{2n+2}$. Moreover, $-n<-\frac{n(2n+1)}{2n+2}<0,\,h(-n)=h(0)=0<1$, $h(1)=1+n>1$ and $h(-n-1)=(n+1)^{2n+1}>1$. Therefore $h(t)-1$ has exactly two real roots $a_{+}$ and $a_{-}$ with
$a_{+}\in(0,1),\qquad a_{-}\in(-n-1,-n).$
Moreover, by examining the signs of $h^{\prime}(a_{+})$ and $h^{\prime}(a_{-})$, we see that they are both simple roots. Noting that the critical points of $f$ are of the form $(t^{n},\ldots,t^{n},t^{n+1})$, we have
Proposition 3.2.
The conifold point of $f$ is $\mathbf{x}_{\rm con}:=(a_{+}^{n},...,a_{+}^{n},a_{+}^{n+1})$, and $g(a_{+})$ is the critical value $T_{\rm con}$ of $f$ at $\mathbf{x}_{\rm con}$.
Recall the following Rouché’s theorem in classical complex analysis (see, e.g., [GKR, Theorem 6.24]):
Proposition 3.3 (Rouché’s theorem).
For two functions $f_{1}$ and $f_{2}$ holomorphic in a disk $D_{R}$, if $|f_{2}|<|f_{1}|$ on $\partial D_{R}$, then $f_{1}$ and $f_{1}+f_{2}$ have the same number of zeros (counted with multiplicity) in $D_{R}$.
Proposition 3.4.
In $D_{1}$, $h(t)-1$ has exactly $12n+$ 1 roots with multiplicity counted.
Proof.
We write $h(t)-1=f_{1}(t)+f_{2}(t)$ with $f_{1}(t)=nt^{2n+1}$ and $f_{2}(t)=t^{2n+2}-1$.
When $n>2$, we have
$|f_{2}(t)|=|t^{2n+2}-1|\leq|t|^{2n+2}+1=2<n=|nt^{2n+1}|=|f_{1}(t)|,\quad\forall t\in\partial D_{1}.$
By Rouché’s theorem, $h(t)-1$ has $12n+$ 1 roots in $D_{1}$.
When $n=2$, we have
$|f_{2}(t)|\leq(1+\epsilon)^{6}+1<2(1+\epsilon)^{5}=|f_{1}(t)|,\quad\forall t\in\partial D_{1+\epsilon},$
provided that $0<\epsilon<{2\over 3}$ (where we notice $(1+\epsilon)^{6}+1-2(1+\epsilon)^{5}$ is strictly decreasing in $[0,{2\over 3})$). Therefore $h(t)-1$ has 5 roots in $D_{1}$ by taking $\epsilon\to+0$. ∎
Lemma 3.5.
For any Laurent polynomial $p(t)$ in $\mathbb{R}_{\geq 0}[t,t^{-1}]$, we have
$\max_{t\in\partial D_{R}}|p(t)|=p(R).$
Proof.
Write $p(t)=\sum_{i}u_{i}t^{i}$, where $u_{i}\in\mathbb{R}_{\geq 0}$ for all $i$. We have
$p(t)|_{t=R}\leq\max\limits_{t\in\partial D_{R}}|p(t)|\leq\sum\limits_{i}u_{i}|t|^{i}\Big{|}_{t\in\partial D_{R}}=\sum\limits_{i}u_{i}R^{i}=p(R).$
Therefore the statement follows. ∎
Corollary 3.6.
All the roots of $h(t)-1$ but $a_{-}$ are in the annulus $N_{[a_{+},1]}$.
Proof.
By Lemma 3.5, we have $1=h(a_{+})=\max_{t\in\partial D_{a_{+}}}|h(t)|$. By the maximum modulus principle, $|h(t)|<1$ for any $t\in\mathring{D}_{a_{+}}$. Hence the interior $\mathring{D}_{a_{+}}$ contains no roots of $h(t)-1$. By Proposition 3.4 and noting $a_{-}\in(-n-1,-n)$, the statement follows. ∎
Recall that the functions $g$ and $\tilde{g}$ in (11) take the same value at the roots of $h-1$.
Lemma 3.7.
$\tilde{g}(a_{+})=\max_{t\in N_{[a_{+},1]}}|\tilde{g}(t)|>\tilde{g}(1).$
Proof.
Notice that $\tilde{g}(t)$ is holomorphic on $N_{[a_{+},1]}$ and $\tilde{g}(t)\in\mathbb{R}_{\geq 0}[t,t^{-1}]$. By the maximum modulus principle and Lemma 3.5, we have
$\max_{t\in N_{[a_{+},1]}}|\tilde{g}(t)|=\max_{t\in\partial D_{a_{+}}\cup\partial D_{1}}|\tilde{g}(t)|=\max\{\tilde{g}(a_{+}),\tilde{g}(1)\}=\tilde{g}(a_{+}).$
Here we have $\tilde{g}(1)<\tilde{g}(a_{+})$ by noting $\tilde{g}^{\prime}(x)=(x^{n}+\frac{2}{x^{n+1}})^{\prime}<0$ for any $x\in(0,1)$. ∎
The nonlinear programming Problem 3.1 can be answered by the following proposition.
Proposition 3.8.
Let $n\geq 2$. Take any $\alpha\in\mathbb{C}\setminus\{a_{-},a_{+}\}$ with $h(\alpha)=1$. We have
  1. (1)
    $|g(\alpha)|<g(a_{+})<2n+3$.
  2. (2)
    $n^{n}-1<(-1)^{n}g(a_{-})<n^{n}$.
  3. (3)
    If $n\geq 3$, then $|g(a_{-})|>g(a_{+})$; if $n=2$, $g(a_{+})>|g(a_{-})|$.
Proof.
Recall that the functions $g,\tilde{g},\check{g}$ in (11) take the same value at the roots of $h-1$. It follows from Corollary 3.6 and Lemma 3.7 that $|g(\alpha)|=|\tilde{g}(\alpha)|\leq\tilde{g}(a_{+})$, with equality only if $\alpha\in\partial D_{a_{+}}$. Tracking the inequalities in the proof of Lemma 3.5 with respect to $\check{g}(t)$, we see that $|\check{g}(\alpha)|=|\tilde{g}(\alpha)|=\tilde{g}(a_{+})=\check{g}(a_{+})$ only if $|2\alpha+2n+1|=2|\alpha|+2n+1$, implying $\alpha\in\mathbb{R}_{>0}$. However, this contradicts to the fact that $a_{+}$ is the only positive root of $h(t)-1$. Hence $|g(\alpha)|<g(a_{+})$. Since $a_{+}\in(0,1)$, we have $g(a_{+})=\check{g}(a_{+})<\check{g}(1)=2n+3$ (by noting that $\check{g}$ is strictly increasing in $\mathbb{R}_{>0}$). Hence, statement (1) holds.
Assume $n$ to be odd. Viewing $h,g,\check{g}$ as real functions on $[-n-1,-n]\subset\mathbb{R}$, we have
$0<-\check{g}^{\prime}(x)=-((2n+2)x+2n^{2}+n)x^{n-1}<-((2n+2)x+2n^{2}+n)x^{2n}=-h^{\prime}(x).$
Hence, $g(a_{-})=\check{g}(a_{-})>\check{g}(-n)=-n^{n}$ and $g(a_{-})<\check{g}(-n)+1=-n^{n}+1$, by noting
$\check{g}(a_{-})-\check{g}(-n)=\int_{a_{-}}^{-n}-\check{g}^{\prime}(x)dx<\int_{a_{-}}^{-n}-h^{\prime}(x)dx=h(a_{-})-h(-n)=1-0=1.$
The argument for even $n$ is the same. Thus statement (2) holds.
For $n\geq 3$, $|g(a_{-})|>n^{n}-1>2n+3>g(a_{+})$. For $n=2$, we have
$g(a_{+})=\check{g}(a_{+})>\check{g}(\tfrac{4}{5})=2(\tfrac{4}{5})^{3}+5(\tfrac{4}{5})^{2}>2^{2}>g(a_{-})>2^{2}-1=3.$
Therefore statement (3) holds. ∎
The following proposition serves as the conclusion of this section. Recall $n\geq 1$.
Theorem 3.9.
Conjecture $\mathcal{O}$ holds for $X_{n}$ if and only if either $n$ is even or $n=1$.
Proof.
If $n=1$, then Conjecture $\mathcal{O}$ holds by [HKLY]. Assume $n\geq 2$ now. By Proposition 2.7, Property $\mathcal{O}$ (1) holds if and only if $n$ is even. Notice that $X_{n}$ has Fano index one, and that the set $\{|f(\mathbf{x})|:\mathbf{x}\mbox{ is a critical point of }f\}$ achieves the maximum at $g(t)$ for a unique root $t$ of $h(t)-1$ by Proposition 3.8. Thus Property $\mathcal{O}$ (2) holds as well, provided that Property $\mathcal{O}$ (1) holds. ∎
4. Gamma conjecture I for $X_{n}$
In this section, we will show in Theorem 4.5 that Gamma conjecture I does not hold for $X_{n}=\mathbb{P}_{\mathbb{P}^{n}}({\mathcal{O}}\oplus{\mathcal{O}}(n))$ when $n\geq 4$ is even.
4.1. Gamma class and conifold point
In this section, we follow the notation in Section 2.2, and single out a result implicitly contained in the proof of [GaIr, Theorem 6.3] in terms of Theorem 4.1.
Recall that for an $N$-dimensional toric Fano manifold $X$, its mirror superpotential $f$ is given by (5). The restriction $f|_{({\mathbb{R}}_{>0})^{N}}\colon({\mathbb{R}}_{>0})^{N}\to{\mathbb{R}}$ to the positive-real locus is proper, bounded below and strictly convex in the sense that the matrix $(\partial_{\log x_{i}}\partial_{\log x_{j}}f({\mathbf{x}}))_{i,j}$ is positive definite for every ${\mathbf{x}}\in({\mathbb{R}}_{>0})^{N}$. Therefore it attains a global minimum at a unique point $\mathbf{x}_{\text{con}}\in({\mathbb{R}}_{>0})^{N}$ called the conifold point [Ga1]; ${\mathbf{x}}_{\text{con}}$ is clearly a non-degenerate critical point of $f$.
Theorem 4.1.
There are classes $\alpha_{0},\alpha_{1},\alpha_{2},\dots\in H^{\bullet}(X)$ with $\alpha_{0}\neq 0$ such that we have the asymptotic expansion
$e^{\frac{T_{\rm con}}{z}}S(z)z^{-\mu}z^{c_{1}}\widehat{\Gamma}_{X}\sim\alpha_{0}+\alpha_{1}z+\alpha_{2}z^{2}+\cdots$
as $z\to 0$ along the angular sector $|\arg z|<\pi/2$.
Proof.
It follows from the argument in [Iri09, Section 4.3.1] that
$\langle\phi,S(z)z^{-\mu}z^{c_{1}}\widehat{\Gamma}_{X}\rangle^{X}=z^{-\frac{N}{2}}\int_{({\mathbb{R}}_{>0})^{N}}e^{-{f(x)\over z}}\varphi(x,-z)\frac{dx_{1}\cdots dx_{N}}{x_{1}\cdots x_{N}},$
where $z>0$, $\phi\in H^{\bullet}(X)$, and $\varphi(x,z)\in{\mathbb{C}}[x_{1}^{\pm},...,x_{N}^{\pm},z]$ is such that the class $[\varphi(x,z)\frac{dx_{1}\cdots dx_{N}}{x_{1}\cdots x_{N}}]$ corresponds to $\phi$ under the mirror isomorphism in [Iri09, Proposition 4.8]. In particular, when $\phi=1$, we have $\varphi=1$.
Note that $f|_{({\mathbb{R}}_{>0})^{N}}$ is a real-valued function. So for $z\to 0$ along the sector $|\arg z|<\pi/2$, one can apply the Laplace method (see e.g. [Fed]) to get the asymptotic expansion
$\langle\phi,S(z)z^{-\mu}z^{c_{1}}\widehat{\Gamma}_{X}\rangle^{X}\sim\frac{(2\pi)^{\frac{N}{2}}}{\left|\det\Big{(}\partial_{\log x_{i}}\partial_{\log x_{j}}f(\mathbf{x}_{\text{con}})\Big{)}\right|^{\frac{1}{2}}}e^{-\frac{T_{\rm con}}{z}}\Big{[}\varphi(\mathbf{x}_{\rm con},z)+O(z)\Big{]}.$
This implies that there exist classes $\alpha_{0},\alpha_{1},\alpha_{2},\dots\in H^{\bullet}(X)$ such that
$e^{\frac{T_{\rm con}}{z}}S(z)z^{-\mu}z^{c_{1}}\widehat{\Gamma}_{X}\sim\alpha_{0}+\alpha_{1}z+\alpha_{2}z^{2}+\cdots$
as $z\to 0$ along the angular sector $|\arg z|<\pi/2$. Now we use $\phi=1$ and $\varphi=1$ to get
$\langle 1,\alpha_{0}\rangle^{X}=\lim\limits_{z\to+0}\langle 1,e^{\frac{T_{\rm con}}{z}}S(z)z^{-\mu}z^{c_{1}}\widehat{\Gamma}_{X}\rangle^{X}=\frac{(2\pi)^{\frac{N}{2}}}{\left|\det\Big{(}\partial_{\log x_{i}}\partial_{\log x_{j}}f(\mathbf{x}_{\text{con}})\Big{)}\right|^{\frac{1}{2}}}\neq 0,$
implying that $\alpha_{0}$ is non-zero. ∎
Remark 4.2.
The arguments here are essentially those in [GaIr, proof of Theorem 6.3]. However, the assumption here is weaker, since we do not impose any condition on $T_{\rm con}$.
Theorem 4.3.
The non-zero class $\alpha_{0}\in H^{\bullet}(X)$ in Theorem 4.1 is an eigenvector of $\hat{c}_{1}$ with eigenvalue $T_{\rm con}$.
Proof.
Recall that $s(z):=S(z)z^{-\mu}z^{c_{1}}\widehat{\Gamma}_{X}$ is $\nabla$-flat, i.e.
$z\partial_{z}s(z)=z^{-1}\hat{c}_{1}s(z)-\mu s(z).$
Theorem 4.1 implies there exist $\alpha_{0},\alpha_{1},\alpha_{2},...\in H^{\bullet}(X)$ such that
$e^{T_{\text{con}}/z}s(z)\sim\alpha_{0}+\alpha_{1}z+\alpha_{2}z^{2}+\cdots$
as $z\to 0$ in the angular sector $|\arg z|<\pi/2$. Observe that $\tilde{s}(z)=e^{T_{\text{con}}/z}s(z)$ satisfies the differential equation
$z\partial_{z}\tilde{s}(z)=z^{-1}(\hat{c}_{1}-T_{\text{con}})\tilde{s}(z)-\mu\tilde{s}(z)$
Using the fact that the differentiation commutes with asymptotic expansion, we obtain
$0=(\hat{c}_{1}-T_{\text{con}})\alpha_{0}$
from the leading term. The conclusion follows. ∎
4.2. Gamma conjecture I for $X_{n}$
Following the notation in Section 3, we recall that $T_{\rm con}=g(a_{+})$ for $X_{n}$.
Theorem 4.4.
When $n\in\{1,2\}$, Gamma conjecture I holds for $X_{n}$.
Proof.
Note that $X_{1}$ is the blow-up of ${\mathbb{P}}^{2}$ at one point, which is a del Pezzo surface. This case was proved in [HKLY].
When $n=2$, from Proposition 3.8 (3), we see that $T_{\rm con}$ is the spectral radius of $\hat{c}_{1}$, and it is of multiplicity one. So the statement follows from [GaIr, Theorem 6.3]. ∎
Theorem 4.5.
When $n\geq 3$ is even, Gamma conjecture I does not hold for $X_{n}$.
Proof.
From Proposition 3.8, we see that the spectral radius $\rho$ of $\hat{c}_{1}$, given by $-g(a_{-})$, is strictly larger than $T_{\rm con}$. Note that
$e^{\frac{\rho}{z}}S(z)z^{-\mu}z^{c_{1}}\widehat{\Gamma}_{X_{n}}=e^{\frac{\rho-T_{\rm con}}{z}}\cdot e^{\frac{T_{\rm con}}{z}}S(z)z^{-\mu}z^{c_{1}}\widehat{\Gamma}_{X_{n}}.$
It follows from Theorem 4.1 that $e^{\frac{\rho}{z}}S(z)z^{-\mu}z^{c_{1}}\widehat{\Gamma}_{X_{n}}$ does not have moderate growth as $z\to+0$. Therefore the principal asymptotic class $A_{X_{n}}$ of $X_{n}$ is not given by its Gamma class. ∎
Remark 4.6.
We shall determine $A_{X_{n}}$ in Section 7.4.
When $n\geq 3$ is odd, Conjecture ${\mathcal{O}}$ does not hold for $X_{n}$ (Proposition 3.9), so we cannot talk about Gamma conjecture I for $X_{n}$. However, Proposition 3.8 shows that $X_{n}$ satisfies condition (*1.3), i.e. $\rho^{\prime}=\max\{\operatorname{Re}(\lambda):\lambda\in\operatorname{Spec}(\hat{c}_{1})\}$ is a simple eigenvalue of $\hat{c}_{1}$ and that any other eigenvalue has strictly smaller real part. Therefore we can talk about the variant of Gamma conjecture I in Section 2.1.3.
Theorem 4.7.
When $n\geq 3$ is odd, the variant of Gamma conjecture I holds for $X_{n}$.
Proof.
From Proposition 3.8, we see that $\rho^{\prime}=T_{\rm con}$. Now the statement follows from Theorem 4.1. ∎
5. Towards modified Gamma conjecture I
Our Theorems 4.5 and 4.7 indicate that the original statement of Gamma conjecture I has to be revised. In this section, we discuss potential modifications of Gamma conjecture I for general Fano manifolds, keeping the specialization $\mathbf{q}=\mathbf{1}$.
5.1. Numerical analysis for eigenvalues of $\hat{c}_{1}$
We start with an analysis on the eigenvalues of $\hat{c}_{1}$. In light of the example $X_{n}$, the variant of Gamma conjecture I may hold under a condition weaker than condition (a) in Definition 2.8 (which does not hold for $X_{n}$, $n\geq 3$). Therefore we first introduce the Property $\mathcal{O}_{B}$ for toric Fano manifolds as follows, which involves the mirror Landau-Ginzburg model on the B-side. The corresponding Property $\mathcal{O}_{A}$ on the A-side will be discussed in the next section.
Definition 5.1 (Property $\mathcal{O}_{B}$ for toric Fano manifolds).
We say that a toric Fano manifold $X_{\Sigma}$ satisfies the Property ${\mathcal{O}}_{B}$, if its mirror superpotential $f=f_{\Sigma}$ satisfies following two conditions:
  1. (a)
    There is a unique critical point ${\mathbf{x}}$ of $f$ such that $\operatorname{Re}(f({\mathbf{x}}^{\prime}))<\operatorname{Re}(f({\mathbf{x}}))$ for any other critical point ${\mathbf{x}}^{\prime}$ of $f$. Moreover the critical point ${\mathbf{x}}$ is non-degenerate.
  2. (b)
    the criticl point ${\mathbf{x}}$ in (a) is the conifold point ${\mathbf{x}}_{\text{con}}$.
The Property ${\mathcal{O}}_{B}$(a) is equivalent to condition (*1.3) by Proposition 2.7. The same argument as in [GaIr, Theorem 6.3] shows:
Proposition 5.2.
Suppose that a toric Fano manifold $X_{\Sigma}$ satisfies Property ${\mathcal{O}}_{B}$. Then $X_{\Sigma}$ satisfies the variant of Gamma conjecture I.
By numerical computations for mirrors of low-dimensional toric Fano manifolds, we have obtained [^2] Table 1 was obtained by using Mathematica with all versions 13.0, 10.0 and 8.0 involved. Since the numerical computation may contain errors due to the command “NSolve”, this table should be used for reference only. Table 1.
Table 1. Eigenvalue properties of $\hat{c}_{1}$ for low-dimensional toric Fano manifolds
Dimension Number of toric Property ${\mathcal{O}}$(1) Condition (*1.3) Property ${\mathcal{O}}_{B}$
Fano manifolds fails fails fails
1 1 0 0 0
2 5 0 0 0
3 18 0 0 0
4 124 8 0 4
5 866 111 9 44
6 7622 1409 245 626
As we can see from Table 1, most of the toric Fano manifolds $X_{\Sigma}$ of dimension $\leq 6$ satisfy Property $\mathcal{O}(1)$ and Property $\mathcal{O}_{B}$. Almost all of them statisfy condition (*1.3) (equivalent to Property $\mathcal{O}_{B}$(a)).
Property $\mathcal{O}_{B}$ for general Fano manifolds $X$ requires a priori a construction of the mirror Landau-Ginzburg model $(\check{X},f)$. Whenever $X$ admits a toric degeneration to toric Fano variety $X_{0}$, we may use the superpotential $f_{0}$ mirror to $X_{0}$, and then talk about Property $\mathcal{O}_{B}$ with respect to $f$ and $f_{0}=f_{\Sigma}$. Note however that the toric mirror $f_{0}$ may miss some critical points.
5.2. Discussions on modifications of Gamma conjecture I
For the Gamma class of a Fano manifold, we propose the following conjecture.
Conjecture 5.3 (weak Gamma conjecture I).
For a Fano manifold $X$, there exist classes $\alpha_{0},\alpha_{1},\alpha_{2},\dots\in H^{\bullet}(X)$ such that $\alpha_{0}$ is a non-zero eigenvector of $\hat{c}_{1}$ with eigenvalue $T>0$, and
$e^{T/z}S(z)z^{-\mu}z^{c_{1}}\widehat{\Gamma}_{X}\sim\alpha_{0}+\alpha_{1}z+\alpha_{2}z^{2}+\cdots$
as $z\to+0$.
This is a weak version of Gamma conjecture I. One can verify that if such an infinite tuple $(T,\alpha_{0},\alpha_{1},...)$ exists, then it is unique.
We collect some evidence for this conjecture. If $X$ satisfies Conjecture ${\mathcal{O}}$ and Gamma conjecture I, then it follows from [GGI, Proposition 3.3.1] that such an expansion exists with $T$ being the spectral radius of $\hat{c}_{1}$. If $X$ is a toric Fano manifold, then Theorems 4.1 and 4.3 imply that such an expansion holds with $T=T_{\rm con}$.
From the viewpoint of mirror symmetry together with the evidence for Conjecture 5.3 (e.g. Theorem 4.3), there should be the notion of Property $\mathcal{O}_{A}$, the incomplete statement of which is given as follows.
Pre-Definition 5.4 (Property $\mathcal{O}_{A}$).
We say that a Fano manifold $X$ satisfies Property ${\mathcal{O}}_{A}$ if the following holds:
  1. (a)
    $X$ satisfies condition (*1.3), i.e. $\hat{c}_{1}$ has a simple rightmost eigenvalue; in this case the rightmost eigenvalue is necessarily $\rho^{\prime}=\max\{\operatorname{Re}(\lambda):\lambda\in\operatorname{Spec}(\hat{c}_{1})\}$;
  2. (b)
    $X$ satisfies a as-yet-unidentified hypothesis (see below for elaboration) on the eigenvector of $\hat{c}_{1}$ with eigenvalue $\rho^{\prime}$, which corresponds to the conifold point condition in the case of toric Fano manifolds.
In particular for a toric Fano manifold $X$, the expecting Property ${\mathcal{O}}_{A}$ should be equivalent to Property ${\mathcal{O}}_{B}$. In view of Proposition 5.2, we can expect the following formulation, as a modification of the original Gamma conjecture I.
Pre-Conjecture 5.5 (Modified Gamma conjecture I).
Suppose that a Fano manifold $X$ satisfies Property ${\mathcal{O}}_{A}$. Then $X$ satisfies the variant of Gamma conjecture I, i.e the principal asymptotic class $A_{X}$ is given by the Gamma class $\widehat{\Gamma}_{X}$.
Here we propose two candidates of the expecting Property $\mathcal{O}_{A}$(b), inspired from some (numerically) computational examples as well as the Perron-Frobenius theory of non-negative matrices.
(cb1) The accompanying (companion) matrix of the characteristic polynomial of $\hat{c}_{1}$ admits a unique non-negative eigenvector $\mathbf{v}_{+}$; moreover, $\rho^{\prime}$ is the corresponding eigenvalue of $\mathbf{v}_{+}$.
(cb2) Take a basis $\{\phi_{i}\}$ of $H^{\bullet}(X)$ such that the associated matrix of the operator $c_{1}(X)\cup$ is non-negative. The associated matrix of $\hat{c}_{1}$ with respect to the basis $\{\phi_{i}\}$ admits a unique non-negative eigenvector $\mathbf{v}_{+}$; moreover, $\rho^{\prime}$ is the corresponding eigenvalue of $\mathbf{v}_{+}$.
We warn that the existence and the uniqueness in the candidates are not guaranteed yet. For (cb2), the existence of the expected basis $\{\phi_{i}\}$ could be ensured by the hard Lefschetz $\mathfrak{sl}_{2}$ action on $H^{*}(X)$, but we still need to verify that the formulation is independent of the choices of such bases. Apart form this, (cb2) holds for any flag variety $G/P$ and toric Fano surfaces. By numerical computations, (cb1) holds for all del Pezzo surfaces. We would like to point out that the formulation of Property $\mathcal{O}_{B}$(b) is not quite intrinsic yet. We need to have certain positivity for the superpotential $f$, in order to define the conifold point in general. The candidates are proposed in attempt to match the positivity. We hope to explore this more carefully in the future.
Property ${\mathcal{O}}_{B}$
B-analogue of
Property ${\mathcal{O}}$
Variant of
Gamma Conjecture I
Condition (*1.3)
= Property ${\mathcal{O}}_{B}$(a)
= Property ${\mathcal{O}}_{A}$(a)
underlyingassumption
as-yet-unidentified
Property ${\mathcal{O}}_{A}$
?no relation
Property ${\mathcal{O}}$(1)
Figure 1. A logical relationship between various conditions.
6. Gamma conjecture I over the Kähler moduli space
Recall that the small quantum product $\star$ is parametrized by ${\mathbf{q}}$, which lies in the “Kähler moduli space” $H^{2}(X,{\mathbb{Z}})\otimes{\mathbb{C}}^{\times}$. While both Conjecture ${\mathcal{O}}$ and Gamma conjecture I focus on the quantum multiplication at ${\mathbf{q}}=\mathbf{1}$, no theoretical justification (beyond empirical evidence) existed for this specific specialization. In this section, we explore these conjectures for general values of ${\mathbf{q}}$.
We will also demonstrate that varying ${\mathbf{q}}$ is closely related to birational geometry of the Fano manifold. When ${\mathbf{q}}$ varies and approaches a boundary point of the Kähler moduli space, the spectrum of quantum cohomology assembles in a specific way, reflecting the geometry of the corresponding extremal contraction; such phenomena have been observed in [GoWo, Iri20b, IrKo, Iri23b]. In the next Section 7, we shall investigate this phenomenon in more details for the example $X_{n}={\mathbb{P}}_{{\mathbb{P}}^{n}}({\mathcal{O}}\oplus{\mathcal{O}}(n))$.
6.1. Quantum cohomology and extremal contractions
As mentioned above, the small quantum product $\star$ of a Fano manifold $X$ is parametrized by a point ${\mathbf{q}}$ in the Kähler moduli space
$\mathcal{M}_{X}:=\operatorname{Hom}(H_{2}(X,{\mathbb{Z}})/{\text{torsion}},{\mathbb{C}}^{\times})\cong H^{2}(X,{\mathbb{Z}})\otimes{\mathbb{C}}^{\times}$
where the monomial ${\mathbf{q}}^{\mathbf{d}}$ in the quantum product $\star$ is identified with the function $\mathcal{M}_{X}\to{\mathbb{C}}^{\times}$ given by the evaluation at ${\mathbf{d}}\in H_{2}(X,{\mathbb{Z}})$. We also introduce a linear coordinate $\tau\in H^{2}(X,{\mathbb{C}})$ on $\mathcal{M}_{X}$ related to ${\mathbf{q}}$ via
${\mathbf{q}}^{\mathbf{d}}=e^{\tau\cdot{\mathbf{d}}}$
for ${\mathbf{d}}\in H_{2}(X,{\mathbb{Z}})$. The space $H^{2}(X,{\mathbb{C}})=H^{1,1}(X)$ of $\tau$ is identified with the universal covering of $\mathcal{M}_{X}$. Let $\operatorname{NE_{\mathbb{N}}}(X)\subset H_{2}(X,{\mathbb{Z}})/{\text{torsion}}$ denote the monoid generated by effective curves. This monoid partially compactifies $\mathcal{M}_{X}$:
$\mathcal{M}_{X}\hookrightarrow\overline{\mathcal{M}}_{X}:=\operatorname{Hom}(\operatorname{NE_{\mathbb{N}}}(X),{\mathbb{C}}^{\times})$
where the coordinate ring of ¯$\overline{\mathcal{M}}_{X}$ is the monoid ring ${\mathbb{C}}[\operatorname{NE_{\mathbb{N}}}(X)]$. The small quantum product $\star$ extends to a family of supercommutative product structures over ¯$\overline{\mathcal{M}}_{X}$: we call ¯$\overline{\mathcal{M}}_{X}$ the partially compactified Kähler moduli space.
Let $\operatorname{\overline{NE}}(X)\subset H_{2}(X,{\mathbb{R}})$ be the closure of the cone generated by $\operatorname{NE_{\mathbb{N}}}(X)$; it is called the Mori cone. An extremal ray $R={\mathbb{R}}_{\geq 0}{\mathbf{d}}$ is a one-dimensional face of the cone $\operatorname{\overline{NE}}(X)$ generated by a class ${\mathbf{d}}$ with $c_{1}(X)\cdot{\mathbf{d}}>0$. By the Contraction Theorem [KMM, Theorem 3.1.2], an extremal ray $R$ associates a contraction $f\colon X\to Y$ to a normal projective variety $Y$ such that a curve in $X$ contracts to a point in $Y$ if and only if its homology class lies in $R$. We call $f$ an extremal contraction. The extremal ray $R$ also defines a ring homomorphism
${\mathbb{C}}[\operatorname{NE_{\mathbb{N}}}(X)]\to{\mathbb{C}}[R_{\mathbb{N}}],\quad{\mathbf{q}}^{\mathbf{d}}\mapsto\begin{cases}{\mathbf{q}}^{\mathbf{d}}&\text{if ${\mathbf{d}}\in R$}\\ 0&\text{otherwise}\end{cases}$
where $R_{\mathbb{N}}=R\cap\operatorname{NE_{\mathbb{N}}}(X)$. This defines a line
$\mathcal{C}_{f}:=\operatorname{Spec}{\mathbb{C}}[R_{\mathbb{N}}]\subset\overline{\mathcal{M}}_{X}$
in the boundary of the partially compactified Kähler moduli space. We call $\mathcal{C}_{f}$ the $f$-boundary of ¯$\overline{\mathcal{M}}_{X}$.
We want to consider the limit where the parameter ${\mathbf{q}}\in\mathcal{M}_{X}$ approaches $\mathcal{C}_{f}$. Given an ample class $\omega_{Y}$ on the base $Y$ of the contraction, we can consider the flow $s\mapsto{\mathbf{q}}_{s}$ on $\mathcal{M}_{X}$ given by
${\mathbf{q}}_{s}=e^{-sf^{*}\omega_{Y}}{\mathbf{q}}\qquad\text{(or more precisely, ${\mathbf{q}}_{s}^{\mathbf{d}}=e^{-sf^{*}\omega_{Y}\cdot{\mathbf{d}}}{\mathbf{q}}^{\mathbf{d}}$).}$.
As $s\to\infty$, the flow ${\mathbf{q}}_{s}$ converges to a point in the $f$-boundary $\mathcal{C}_{f}$. In terms of the linear coordinate $\tau\in H^{2}(X,{\mathbb{C}})$, the same flow is given by
$s\mapsto\tau_{s}=\tau-sf^{*}\omega_{Y}.$
Under the $s\to\infty$ limit, the small quantum product of $X$ at ${\mathbf{q}}_{s}$ approaches the $f$-exceptional quantum product $\star_{\text{exc}}$ defined by
$\alpha\star_{\text{exc}}\beta=\sum_{i}\sum_{{\mathbf{d}}\in R_{\mathbb{N}}}\langle\alpha,\beta,\phi^{i}\rangle_{\mathbf{d}}{\mathbf{q}}^{\mathbf{d}}\phi_{i}.$
This defines a family of supercommutative rings over $\mathcal{C}_{f}$. A similar product has been defined by Ruan [Rua06] in connection with his famous crepant resolution conjecture. The $f$-exceptional quantum product, which is easier to compute, leads to a decomposition of the original quantum cohomology. When ${\mathbf{q}}$ is sufficiently close to a boundary point $\overline{{\mathbf{q}}}\in\mathcal{C}_{f}={\mathbb{C}}$, the quantum cohomology ring $QH(X)|_{\mathbf{q}}=(H^{*}(X),\star_{\mathbf{q}})$ decomposes into subrings indexed by maximal ideals $\mathfrak{m}$ of $(H^{\text{ev}}(X),\star_{\text{exc},\overline{{\mathbf{q}}}})$ (i.e. points of the 00-dimensional scheme $\operatorname{Spec}(H^{\text{ev}}(X),\star_{\text{exc},\overline{{\mathbf{q}}}})$) for the non-zero complex number $\overline{{\mathbf{q}}}$:
(12) $QH(X)\bigr{|}_{{\mathbf{q}}}\cong\bigoplus_{\mathfrak{m}}A_{\mathfrak{m}}.$
Question 6.1.
Can we identify each factor $A_{\mathfrak{m}}$ with quantum cohomology of another space in a natural way?
Examples (see below) show that $A_{\mathfrak{m}}$ can arise as the (big) quantum cohomology of the base space $Y$, the flip $X^{\dagger}\to Y$ of $f$, or a space related to the exceptional set of $f$.
Remark 6.2.
The $f$-boundary $\mathcal{C}_{f}$ and $f$-exceptional quantum product $\star_{\text{exc}}$ can be defined beyond the Fano case. For the degree reason, $\star_{\text{exc}}$ is always convergent (polynomial) even if the original quantum product $\star$ is not known to be convergent. In general, the decomposition (12) makes sense for ${\mathbf{q}}$ in the formal neighbourhood of $\overline{{\mathbf{q}}}\in\mathcal{C}_{f}$.
Example 6.3 ([GoWo, Theorem 5.5],[Iri20b, Theorem 1.1]).
Let $X$ be a smooth toric Deligne-Mumford (DM) stack given as a GIT quotient of a vector space and let $X\dasharrow X^{\dagger}$ be a flip (or a discrepant transformation) arising from variation of GIT quotients. Then $QH(X)$ contains $QH(X^{\dagger})$ as a direct summand, possibly after deforming $QH(X^{\dagger})$ to big quantum cohomology.
Example 6.4 ([IrKo, Theorem 5.1]).
Let $X={\mathbb{P}}(V)$ be the projectivization of a vector bundle $V$ of rank $r$ over a smooth base $Y$. The ${\mathbb{P}}^{r-1}$-fibration $X\to Y$ gives rise to a decomposition $QH(X)\cong\bigoplus_{i=1}^{r}QH(Y)_{\sigma_{i}}$, where $QH(Y)_{\sigma_{i}}$ stands for the big quantum cohomology with parameter $\sigma_{i}\in H^{*}(Y)$. This can be viewed as a quantum cohomology analogue of the Leray-Hirsch theorem.
Example 6.5 ([Iri23b, Theorem 1.1]).
Let $\widetilde{X}$ be the blowup of a smooth projective variety $X$ along a smooth centre $Z$ and let $r$ be the codimension of $Z$ in $X$. The blowdown morphism $\widetilde{X}\to X$ gives rise to a decomposition $QH(\widetilde{X})\cong QH(X)_{\tau}\oplus\bigoplus_{i=1}^{r-1}QH(Z)_{\sigma_{i}}$ where $\tau\in H^{*}(X)$, $\sigma_{i}\in H^{*}(Z)$ stand for the big quantum cohomology parameters.
6.2. Principal asymptotic class for general ${\mathbf{q}}$
We discuss the principal asymptotic class for general ${\mathbf{q}}\in\mathcal{M}_{X}=H^{2}(X,{\mathbb{Z}})\otimes{\mathbb{C}}^{\times}$ for a Fano manifold $X$. The discussion here is parallel to Section 2.1.3.
Definition 6.6.
Let ${\mathbf{q}}\in\mathcal{M}_{X}$. We say that $(c_{1}(X)\star_{\mathbf{q}})$ has a simple rightmost eigenvalue $u\in{\mathbb{C}}$ if
  • (1)
    $u$ is a simple root of the characteristic polynomial of $(c_{1}(X)\star_{\mathbf{q}})$ and
  • (2)
    any other eigenvalues $u^{\prime}$ of $(c_{1}(X)\star_{\mathbf{q}})$ satisfy $\operatorname{Re}(u^{\prime})<\operatorname{Re}(u)$.
This condition relaxes Property ${\mathcal{O}}$ in several ways: we do not require $u$ to be positive real or have the largest norm among eigenvalues; also we do not require the condition involving the Fano index $i_{X}$. As remarked in [GGI, Remark 3.1.9], it enables us to define the principal asymptotic class.
Remark 6.7.
(1) When ${\mathbf{q}}$ is real, eigenvalues of $(c_{1}(X)\star_{\mathbf{q}})$ are invariant under complex conjugation and a simple rightmost eigenvalue $u$ (if any) is necessarily real. In particular when $\mathbf{q}=\mathbf{1}$, the notion of a simple rightmost eigenvalue coincides with condition (*1.3) (or Property $\mathcal{O}_{A}$(a)).
(2) For every ${\mathbf{q}}\in\mathcal{M}_{X}$, the set of eigenvalues of the small quantum product $(c_{1}(X)\star_{\mathbf{q}})$ is invariant under mutliplication by $i_{X}$th roots of unity. This is because the complex Euler flow ${\mathbb{C}}\ni s\mapsto{\mathbf{q}}_{s}=e^{sc_{1}(X)}{\mathbf{q}}$ dilates the spectrum of $(c_{1}(X)\star_{\mathbf{q}})$ by $e^{s}$ and ${\mathbf{q}}_{s}={\mathbf{q}}$ for $s=2\pi{\mathbf{i}}/i_{X}$. This implies that the argument of a non-zero rightmost eigenvalue lies in the interval $[-\pi/i_{X},\pi/i_{X}].$
We consider the quantum connection
(13) $\nabla_{z\partial_{z}}\bigr{|}_{\mathbf{q}}=z\partial_{z}-\frac{1}{z}(c_{1}(X)\star_{\mathbf{q}})+\mu$
on the trivial $H^{*}(X)$-bundle over ${\mathbb{P}}^{1}$. This has a fundamental solution of the form (due to Dubrovin; see [GGI, Remark 2.3.2] and references therein)
$S(\tau,z)z^{-\mu}z^{c_{1}(X)}$
where $\tau\in H^{2}(X,{\mathbb{C}})$ is a lift of ${\mathbf{q}}$ (i.e. ${\mathbf{q}}=e^{\tau}$) and $S(\tau,z)$ is a $GL(H^{*}(X))$-valued holomorphic function of $(\tau,z)\in H^{2}(X,{\mathbb{C}})\times({\mathbb{P}}^{1}\setminus\{0\})$. The map $\alpha\mapsto S(\tau,z)z^{-\mu}z^{c_{1}(X)}\alpha$ identifies $H^{*}(X)$ with the space of multi-valued flat sections for $\nabla_{z\partial_{z}}|_{\mathbf{q}}$. This fundamental solution is characterized by the condition that
  • it is also flat with respect to the quantum connection in the $\tau$-direction and;
  • $S(0,z)$ coincides with $S(z)$ in Section 2.1.3.
More precisely, in terms of descendant Gromov-Witten invariants, $S(\tau,z)$ is given by
$S(\tau,z)\alpha=e^{-\tau/z}\alpha+\sum_{i}\sum_{{\mathbf{d}}\neq 0}\left\langle\frac{e^{-\tau/z}\alpha}{-z-\psi},\phi^{i}\right\rangle_{0,2,{\mathbf{d}}}e^{\tau\cdot{\mathbf{d}}}\phi_{i}.$
The following result follows by the same argument as in [GGI, Proposition 3.3.1].
Proposition/Definition 6.8.
Suppose that $(c_{1}(X)\star_{\mathbf{q}})$ has a simple rightmost eigenvalue $u$. Then the space
$\{\alpha\in H^{*}(X):\text{$\|e^{u/z}S(\tau,z)z^{-\mu}z^{c_{1}(X)}\alpha\|=O(z^{-m})$ for some $m\in\mathbb{N}$ as $z\to+0$}\}$
is one-dimensional, where $\tau\in H^{2}(X,{\mathbb{C}})$ is a lift of ${\mathbf{q}}$. The principal asymptotic class $A_{X}(\tau)$ at $\tau$ is a generator of this vector space defined up to constant. Moreover the limit
$\lim_{z\to+0}e^{u/z}S(\tau,z)z^{-\mu}z^{c_{1}(X)}A_{X}(\tau)$
exists and is a $u$-eigenvector of $(c_{1}(X)\star_{\mathbf{q}})$.
Remark 6.9.
The class $A_{X}(\tau)$ depends on the choice of a lift $\tau$ of ${\mathbf{q}}$. We have $A_{X}(\tau+2\pi{\mathbf{i}}\xi)=e^{2\pi{\mathbf{i}}\xi}A_{X}(\tau)$ for $\xi\in H^{2}(X,{\mathbb{Z}})$ since $S(\tau+2\pi{\mathbf{i}}\xi,z)=S(\tau,z)e^{-2\pi{\mathbf{i}}\xi/z}$.
The principal asymptotic class also arises as a limit of the $J$-function. The restriction of the $J$-function (3) to $\tau\in H^{2}(X)$ is given by
$J_{X}(\tau,z)$ $=S(\tau,z)^{-1}1=e^{\tau/z}\left(1+\sum_{i}\sum_{{\mathbf{d}}\neq 0}\left\langle\frac{\phi^{i}}{z(z-\psi)}\right\rangle_{0,1,{\mathbf{d}}}e^{\tau\cdot{\mathbf{d}}}\phi_{i}\right).$
Proposition 6.10.
Suppose that $(c_{1}(X)\star_{\mathbf{q}})$ has a simple rightmost eigenvalue $u$ for ${\mathbf{q}}=e^{\tau}\in\mathcal{M}_{X}$. Then we have the asymptotic expansion
$J_{X}(\tau+c_{1}(X)\log t,1)\sim Ct^{-\frac{\dim X}{2}}e^{ut}(A_{X}(\tau)+\alpha_{1}t^{-1}+\alpha_{2}t^{-2}+\cdots)$
as $t\to+\infty$ on the positive real line, where $C\neq 0$ is a constant and $\alpha_{i}\in H^{*}(X)$. In particular, we have $\lim_{t\to+\infty}[J(\tau+c_{1}(X)\log t,1)]=[A_{X}(\tau)]$ in ${\mathbb{P}}(H^{*}(X))$.
Proof.
This follows by the same argument as the $\tau=0$ case in [GaIr, Proposition 3.8], which is based on [GGI, Propositions 3.6.2, 3.2.1]. The argument there shows that
$z^{-c_{1}(X)}z^{\mu}J_{X}(\tau,z)\sim Ce^{u/z}(A_{X}(\tau)+\alpha_{1}z+\alpha_{2}z^{2}+\cdots)$
as $z\to+0$. The conclusion follows from the identity
$t^{\frac{\dim X}{2}}J_{X}(\tau+c_{1}(X)\log t,1)=z^{-c_{1}(X)}z^{\mu}J_{X}(\tau,z)$
with $t=z^{-1}$. ∎
Corollary 6.11.
The principal asymptotic class $A_{X}(\tau)$, when it is defined, depends only on the class $[\tau]$ in $H^{2}(X,{\mathbb{C}})/{\mathbb{R}}c_{1}(X)$ up to a constant multiple. If $\tau\in H^{2}(X,{\mathbb{R}})$, $A_{X}(\tau)$ is a complex multiple of a real cohomology class.
Gamma conjecture I states that $[A_{X}(0)]=[\widehat{\Gamma}_{X}]$ in ${\mathbb{P}}(H^{*}(X))$. In Theorem 4.5, we show that $[A_{X_{n}}(0)]\neq[\widehat{\Gamma}_{X_{n}}]$ for even $n\geq 4$. This motivates the following question:
Question 6.12.
Does there exist a point ${\mathbf{q}}=e^{\tau}\in\mathcal{M}_{X}$ such that $(c_{1}(X)\star_{\mathbf{q}})$ has a simple rightmost eigenvalue and that $[A_{X}(\tau)]=[\widehat{\Gamma}_{X}]$? Can we find such a $\tau$ within $H^{2}(X,{\mathbb{R}})$? Can we characterize the region in $H^{2}(X)$ where $[A_{X}(\tau)]=[\widehat{\Gamma}_{X}]$ holds?
Even if $[A_{X}(\tau)]\neq[\widehat{\Gamma}_{X}]$, $A_{X}(\tau)$ should be a special class: we can conjecture that $A_{X}(\tau)$ comes from an exceptional object in the derived category. In light of Dubrovin’s conjecture [Du1, Conjecture 4.2.2] and Gamma conjecture II [GGI], we ask the following:
Question 6.13.
When the principal asymptotic class $A_{X}(\tau)$ is defined, does there exist a $K$-class $V\in K^{0}(X)$ of topological vector bundles such that $[A_{X}(\tau)]=[\widehat{\Gamma}_{X}\operatorname{Ch}(V)]$? Here $\operatorname{Ch}(V)=\sum_{k\geq 0}(2\pi{\mathbf{i}})^{k}\operatorname{ch}_{k}(V)$ is the $(2\pi{\mathbf{i}})$-modified Chern character. Does $V$ satisfy $\chi(V,V)=1$? Does it come from an exceptional object in the derived category of coherent sheaves?
If the answer to this question is affirmative at a real $\tau\in H^{2}(X,{\mathbb{R}})$, then Corollary 6.11 implies that $\operatorname{ch}(V)$ lies either in $\bigoplus_{k\geq 0}H^{4k}(X)$ or in $\bigoplus_{k\geq 0}H^{4k+2}(X)$ (i.e. $\operatorname{Ch}(V)$ is real or purely imaginary).
Remark 6.14.
Here we focused on the principal asymptotic class. However, our scope can be broadened to consider (1) higher asymptotic classes associated with a simple eigenvalue of $(c_{1}(X)\star_{\mathbf{q}})$ (see Gamma conjecture II of [GGI]), or even (2) asymptotic subspaces associated with a not necessarily simple eigenvalue of $(c_{1}(X)\star_{\mathbf{q}})$ (coined as “Gamma conjecture III” by the authors of [GGI], see [SaSh, Iri23a]).
6.3. Semisimple case
In this section, we briefly discuss the case where the (big) quantum cohomology is generically semisimple. Note that toric Fano manifolds have generically semisimple (small and big) quantum cohomology.
When the big quantum cohomology is convergent at $\tau\in H^{*}(X)$, we can similarly define the principal asymptotic class $A_{X}(\tau)$. For this we need to replace $(c_{1}(X)\star_{\mathbf{q}})$ in the quantum connection (13) with the Euler multiplication $(E\star_{\tau})$ (we have $E=c_{1}(X)$ for $\tau\in H^{2}(X)$) and use the fundamental solution $S(\tau,z)$ extended to $\tau\in H^{*}(X)$.
Suppose that the big quantum cohomology of $X$ is generically semisimple. Around a semisimple point $\tau_{0}\in H^{*}(X)$, the eigenvalues $u_{1},\dots,u_{s}$ of $(E\star_{\tau})$ form a coordinate system in a neighbourhood of $\tau=\tau_{0}$ [Dub98, Lecture 3]. We can move the eigenvalues $u_{1},\dots,u_{s}$ as long as we can solve the corresponding Riemann-Hilbert problem. In fact, the big quantum cohomology Frobenius manifold can be extended to an open dense subset333the complement of an analytic hypersurface of the universal covering $({\mathbb{C}}^{s}\setminus\text{(big diagonal)})^{~}$ of the configuration space of $s$ points $u_{1},\dots,u_{s}$ [Dub98, Theorem 4.7]. On the semisimple locus, the condition that $(E\star_{\tau})$ has a simple rightmost eigenvalue fails along real-codimension one walls and the big quantum cohomology parameter space is divided into chambers. The principal asymptotic class $A_{X}(\tau)$ stays constant in a chamber, and the dominant asymptotic classes alternate when $\tau$ crosses a wall. Note that this is not a mutation we discuss in §7.4 below.
When the big quantum product is semisimple for generic $\tau\in H^{*}(X,{\mathbb{C}})$, the same holds also for generic real parameters $\tau\in H^{*}(X,{\mathbb{R}})$; the Euler multiplication $(E\star_{\tau})$ has a mutually distinct simple eigenvalues away from a real-codimension one hypersurface in $H^{*}(X,{\mathbb{R}})$. On the real locus $H^{*}(X,{\mathbb{R}})$, we observe two types of open chambers. At a generic $\tau\in H^{*}(X,{\mathbb{R}})$, we have either
  • (i)
    $(E\star_{\tau})$ has a simple rightmost eigenvalue (which is automatically real), or
  • (ii)
    $(E\star_{\tau})$ has a complex-conjugate pair of two simple eigenvalues $u,\overline{u}$ such that all other eigenvalues $u^{\prime}$ satisfy $\operatorname{Re}(u^{\prime})<\operatorname{Re}(u)$.
The locus where both conditions fail is of real-codimension 1 or greater. In fact, if, for some $\tau\in H^{*}(X,{\mathbb{R}})$, $(E\star_{\tau})$ has only simple eigenvalues $u_{1},\dots,u_{s}$ and these conditions fail, then either (a) there are three eigenvalues $u_{1},u_{2},u_{3}$ such that $u_{1}=\overline{u}_{1}$, $u_{3}=\overline{u}_{2}$ and $\operatorname{Re}(u_{1})=\operatorname{Re}(u_{2})=\operatorname{Re}(u_{3})$ or (b) there are four eigenvalues $u_{1},u_{2},u_{3},u_{4}$ such that $u_{2}=\overline{u}_{1}$, $u_{4}=\overline{u}_{3}$ and $\operatorname{Re}(u_{1})=\operatorname{Re}(u_{2})=\operatorname{Re}(u_{3})=\operatorname{Re}(u_{4})$. Such a point $\tau$ belongs to the real-codimension one locus given by the intersection of $H^{*}(X,{\mathbb{R}})$ with the complex hypersurface (a) $u_{1}=(u_{2}+u_{3})/2$ or (b) $u_{1}+u_{2}=u_{3}+u_{4}$. When $\tau\in H^{*}(X,{\mathbb{R}})$ lies in a chamber of type-(ii) with $u=x+{\mathbf{i}}y$, we expect an oscillating asymptotics
$z^{-c_{1}}z^{\mu}J_{X}(\tau,z)$ $\sim e^{x/z}(\operatorname{Re}(e^{iy/z}A)+O(z))$
$=e^{x/z}(\cos(y/z)A_{1}-\sin(y/z)A_{2}+O(z))$
as $z\to+0$, for some $A=A_{1}+{\mathbf{i}}A_{2}\in H^{*}(X,{\mathbb{C}})$. If $\tau\in H^{2}(X,{\mathbb{R}})$, the left hand side equals $t^{\dim X/2}J_{X}(\tau+c_{1}(X)\log t,1)$ with $t=1/z$. The class $A$ here can be viewed as a substitute of the principal asymptotic class in a type-(ii) chamber.
Example 6.15.
Let $f\colon\widetilde{X}\to X$ be a blowup of $X$ at a point. If $\dim X\geq 4$ and a positive real parameter ${\mathbf{q}}=e^{\tau}\in\mathcal{M}_{\widetilde{X}}$ is close to the $f$-boundary, then the point $\tau\in H^{2}(\widetilde{X},{\mathbb{R}})$ lies in a type-(ii) chamber. See Remark 7.2.
7. Gamma conjecture I for $X_{n}$ over the Kähler moduli space
We examine the counterexamples $X_{n}={\mathbb{P}}_{{\mathbb{P}}^{n}}({\mathcal{O}}\oplus{\mathcal{O}}(n))$ to Conjecture ${\mathcal{O}}$/Gamma conjecture I from a birational geometry perspective. The space $X_{n}$ has two extremal contractions: the divisorial contraction $\varphi\colon X_{n}\to{\mathbb{P}}(1^{n+1},n)$ to the weighted projective space and the ${\mathbb{P}}^{1}$-fibration $\pi\colon X_{n}\to{\mathbb{P}}^{n}$. Each contraction corresponds to an extremal ray of the Mori cone and associates a line in the boundary of the Kähler moduli space. We observe that the principal asymptotic class equals the Gamma class $\widehat{\Gamma}_{X_{n}}$ when ${\mathbf{q}}$ is close to the “$\pi$-boundary”. However, when ${\mathbf{q}}$ is close to the “$\varphi$-boundary” and $n$ is even, it deviates from $\widehat{\Gamma}_{X_{n}}$. We also determine the principal asymptotic class in the latter region (that includes ${\mathbf{q}}=\mathbf{1}$) using monodromy and mutation argument.
7.1. Quantum cohomology of $X_{n}$ over the Kähler moduli space
Let us start with the geometric construction of $X_{n}$. It is a toric variety defined by the weight matrix
(14) $\begin{pmatrix}1&1&\cdots&1&0&-n\\ 0&0&\cdots&0&1&1\end{pmatrix}$
of size $2\times(n+3)$. This matrix defines a (×$({\mathbb{C}}^{\times})^{2}$-action on ${\mathbb{C}}^{n+3}$ and $X_{n}$ arises as a GIT quotient of this action. Columns of the matrix represent the classes of toric divisors $D_{1},\dots,D_{n+1},D_{n+2},D_{n+3}$ of $X_{n}$ in $H^{2}(X_{n},{\mathbb{Z}})\cong\operatorname{Hom}(({\mathbb{C}}^{\times})^{2},{\mathbb{C}}^{\times})$, where $D_{i}$ is the zero-set $\{z_{i}=0\}$ of the $i$th homogeneous coordinate in $X_{n}={\mathbb{C}}^{n+3}/\!/({\mathbb{C}}^{\times})^{2}$. They form the divisor diagram in Figure 2. We have two GIT chambers (for stability conditions): the right chamber (the first quadrant) gives rise to $X_{n}$ and the left chamber (spanned by $D_{n+2},D_{n+3}$) gives rise to the weighted projective space
$Y_{n}:={\mathbb{P}}(1^{n+1},n)={\mathbb{P}}(\underbrace{1,\dots,1}_{\text{$n+1$ times}},n)$
as GIT quotients. The birational map $X_{n}\dasharrow Y_{n}$ is an example of toric discrepant transformations studied in [Iri20b].
[…]
Figure 2. The divisor diagram for $X_{n}$ (for $n=4$) and the two GIT chambers.
The $X_{n}$-chamber (spanned by $D_{1},D_{n+2}$) is identified with the ample cone of $X_{n}$. Its dual cone is the Mori cone $\operatorname{\overline{NE}}(X_{n})$ that has two extremal rays. Let $q_{1}$, $q_{2}$ denote the quantum variables corresponding to the primitive generators of the extremal rays, which are dual to $(D_{1},D_{n+2})$. These extremal rays respectively associate the contractions
$\varphi\colon X_{n}\to\overline{Y}_{n},\quad\pi\colon X_{n}\to{\mathbb{P}}^{n}$
where $\varphi$ is the divisorial contraction (contracting $D_{n+3}$) to the coarse moduli space $\overline{Y}_{n}$ of $Y_{n}$ and $\pi$ is the ${\mathbb{P}}^{1}$-fibration. The map $\varphi$ is given by
$\varphi([z_{1},\dots,z_{n+1},z_{n+2},z_{n+3}])=[z_{1}z_{n+3}^{1/n},\dots,z_{n+1}z_{n+3}^{1/n},z_{n+2}].$
The variable $q_{1}$ represents the class of a line in the exceptional divisor $D_{n+3}\cong{\mathbb{P}}^{n}$ and $q_{2}$ represents the class of a fiber of $\pi\colon X_{n}\to{\mathbb{P}}^{n}$.
The Kähler moduli space $\mathcal{M}_{X_{n}}$ of $X_{n}$ is (×$({\mathbb{C}}^{\times})^{2}$ with coordinates ${\mathbf{q}}=(q_{1},q_{2})$. The partially compactified Kähler moduli space ¯$\overline{\mathcal{M}}_{X_{n}}$ is ${\mathbb{C}}^{2}$. The $\varphi$- and $\pi$-boundaries of ¯$\overline{\mathcal{M}}_{X_{n}}$ are given respectively by
$\mathcal{C}_{\varphi}=\text{$q_{1}$-axis}=\{q_{2}=0\},\qquad\mathcal{C}_{\pi}=\text{$q_{2}$-axis}=\{q_{1}=0\}.$
In Sections 2 and 3, we studied the quantum cohomology of $X_{n}$ using mirror symmetry at the specialization $\mathbf{q}=(1,1)$. Let us consider general $\mathbf{q}$ now. The Landau-Ginzburg mirror is a family of functions parametrized by $(q_{1},q_{2})\in\mathcal{M}_{X_{n}}=({\mathbb{C}}^{\times})^{2}$ and is given by
$f_{q_{1},q_{2}}=x_{1}+\cdots+x_{n+1}+y_{1}+y_{2}$
where $(x_{1},\dots,x_{n+1},y_{1},y_{2})$ is required to satisfy the constraints
$x_{1}\cdots x_{n+1}y_{2}^{-n}=q_{1},\qquad y_{1}y_{2}=q_{2}.$
These constraints can be read off from rows of the weight matrix (14). Choosing $(x_{1},\dots,x_{n},y_{2})$ as independent coordinates, we can rewrite $f_{q_{1},q_{2}}$ as the Laurent polynomial
(15) $f_{q_{1},q_{2}}=x_{1}+\cdots+x_{n}+q_{1}\frac{y_{2}^{n}}{x_{1}\cdots x_{n}}+y_{2}+\frac{q_{2}}{y_{2}}.$
Setting $q_{1}=q_{2}=1$ gives the mirror Landau-Ginzburg model in formula (8). We have a ring isomorphism
$QH(X_{n})\bigr{|}_{q_{1},q_{2}}\cong\operatorname{Jac}(f_{q_{1},q_{2}})\cong{\mathbb{C}}[x_{1},y_{2}]/(x_{1}^{n+1}-q_{1}y_{2}^{n},y_{2}(y_{2}+nx_{1})-q_{2})$
[…]
$\operatorname{Jac}(f_{q_{1},0})$ $\cong{\mathbb{C}}[x_{1},y_{2}]/(x_{1}^{n+1}-q_{1}y_{2}^{n},y_{2}(y_{2}+nx_{1})),$
$\operatorname{Jac}(f_{0,q_{2}})$ $\cong{\mathbb{C}}[x_{1},y_{2}]/(x_{1}^{n+1},y_{2}(y_{2}+nx_{1})-q_{2}).$
They are isomorphic to the $\varphi$- and $\pi$-exceptional quantum cohomology. The spectrum of $\operatorname{Jac}(f_{q_{1},0})$ for $q_{1}\neq 0$ consists of two points (one ‘fat’ point of multiplicity $12n+$ 1 and one reduced point) and the spectrum of $\operatorname{Jac}(f_{0,q_{2}})$ for $q_{2}\neq 0$ consists of two ‘fat’ points of multiplicity $n+1$. In the next section we observe how $\operatorname{Spec}(\operatorname{Jac}(f_{q_{1},q_{2}}))$ — which consists of $22n+$ 2 points for generic $q_{1},q_{2}$ — clusters into these spectra, in terms of critical values.
As discussed in [Iri20b], the family of rings $\operatorname{Jac}(f_{q_{1},q_{2}})$ also contains the big quantum cohomology of $Y_{n}$. To see this, we add another boundary $\{q_{1}=\infty\}$ to ¯$\overline{\mathcal{M}}_{X_{n}}$. Introduce new parameters ${\mathfrak{q}}_{1}=q_{1}^{-1/n}$ and ${\mathfrak{q}}_{2}=q_{1}^{1/n}q_{2}$. Under the change of the variables $y_{2}=q_{2}/w$, the Landau-Ginzburg mirror can be rewritten as
$f_{q_{1},q_{2}}={\mathfrak{f}}_{{\mathfrak{q}}_{1},{\mathfrak{q}}_{2}}:=x_{1}+\cdots+x_{n}+\frac{{\mathfrak{q}}_{2}^{n}}{x_{1}\cdots x_{n}w^{n}}+w+\frac{{\mathfrak{q}}_{1}{\mathfrak{q}}_{2}}{w}.$
Along $\{{\mathfrak{q}}_{1}=0\}=\{q_{1}=\infty\}$, this restricts to the Landau-Ginzburg potential ${\mathfrak{f}}_{0,{\mathfrak{q}}_{2}}$ mirror to the small (orbifold) quantum cohomology of $Y_{n}$, i.e.
$QH(Y_{n})\bigr{|}_{{\mathfrak{q}}_{2}}\cong\operatorname{Jac}({\mathfrak{f}}_{0,{\mathfrak{q}}_{2}}).$
By gluing the two families $\{f_{q_{1},q_{2}}\}$, $\{{\mathfrak{f}}_{{\mathfrak{q}}_{1},{\mathfrak{q}}_{2}}\}$ of Landau-Ginzburg potentials, we obtain a global mirror family over the toric orbifold ¯$\overline{\mathcal{M}}_{X_{n}}\cup\{{\mathfrak{q}}_{1}=0\}$ that is associated with the fan as in Figure 2. See Figure 3. Note that the Jacobi ring $\operatorname{Jac}({\mathfrak{f}}_{0,{\mathfrak{q}}_{2}})$ is of rank $12n+$ 1 as one of the critical points of $f_{q_{1},q_{2}}$ goes to infinity in this limit. Near the boundary locus $\{{\mathfrak{q}}_{1}=0\}$, therefore, the Jacobi ring $\operatorname{Jac}(f_{q_{1},q_{2}})$ contains a $(2n+1)$-dimensional subring as a direct summand. We can identify this subring with the big quantum cohomology deformation of $QH(Y_{n})$, where ${\mathfrak{q}}_{1}$ corresponds approximately to the coordinate dual to the twisted sector $\frac{1}{n}(1,\dots,1)$ (supported at the origin of $[{\mathbb{C}}^{n+1}/\mu_{n}]\subset Y_{n}$ with $\mu_{n}$ denoting the cyclic group of order $n$).
[…]
Figure 3. The partially compactified Kähler moduli space ¯$\overline{\mathcal{M}}_{X_{n}}$ together with the (dashed) infinity line $\{{\mathfrak{q}}_{1}=0\}=\{q_{1}=\infty\}$ which is identified with ¯$\overline{\mathcal{M}}_{Y_{n}}$.
7.2. Tropical study of critical points
We study the limiting behaviour of the quantum cohomology ring of $X_{n}$ (or equivalently, the Jacobi ring $\operatorname{Jac}(f_{q_{1},q_{2}})$) near the boundaries $\mathcal{C}_{\pi}$ and $\mathcal{C}_{\varphi}$. More specifically, we study the behaviour of critical values the Landau-Ginzburg potential $f_{q_{1},q_{2}}$ above. We start by noting that there is a conformal symmetry (given by the Euler vector field):
(16) $x_{1}\to e^{s}x_{1},\quad\dots\quad x_{n}\to e^{s}x_{n},\quad y_{2}\to e^{s}y_{2},\quad q_{1}\to e^{s}q_{1},\quad q_{2}\to e^{2s}q_{2}$
under which the Landau-Ginzburg potential (15) is scaled by $e^{s}$. Hence it suffices to consider a family where $q_{1}$ is fixed and only $q_{2}$ changes (the essential parameter is $q_{2}/q_{1}^{2}$). Let $0<\mathsf{T}\ll 1$ be a small tropical parameter and consider the following family:
$f_{q_{1},q_{2}\mathsf{T}^{\lambda}}(x_{1},\dots,x_{n},y_{2})=x_{1}+\cdots+x_{n}+q_{1}\frac{y_{2}^{n}}{x_{1}\cdots x_{n}}+y_{2}+\frac{q_{2}\mathsf{T}^{\lambda}}{y_{2}}$
where $\lambda\in{\mathbb{R}}$. This Laurent polynomial represents a point close to the $\pi$-boundary when $\lambda<0$ and a point close to the $\varphi$-boundary when $\lambda>0$. Following the method in Section 3.2, we find that critical points of $f_{q_{1},q_{2}\mathsf{T}^{\lambda}}$ are given as
$x_{1}=\cdots=x_{n}=q_{1}^{\frac{1}{n+1}}t^{n},\quad y_{2}=t^{n+1}$
for $t$ satisfying the algebraic equation:
(17) $t^{2n+2}+nq_{1}^{\frac{1}{n+1}}t^{2n+1}-q_{2}\mathsf{T}^{\lambda}=0.$
The corresponding critical value $u$ of $f_{q_{1},q_{2}\mathsf{T}^{\lambda}}$ is
(18) $u=2t^{n+1}+(2n+1)q_{1}^{\frac{1}{n+1}}t^{n}.$
[…]
$\min\{(2n+2)e,(2n+1)e,\lambda\}$
is attained by more than one terms. We find that such an $e$ is given as
$e=\begin{cases}0\ \text{ or }\ \lambda/(2n+1)&\text{for $\lambda\geq 0$;}\\ \lambda/(2n+2)&\text{for $\lambda\leq 0$.}\end{cases}$
This shows a bifurcation of critical points at $\lambda=0$: see Figure 4.
$\frac{\lambda}{2}$$\frac{n}{2n+1}\lambda$$\lambda$exponent of $\mathsf{T}$near $\pi$-boundarynear $\varphi$-boundary
Figure 4. Bifurcation of the leading exponent (valuation) of the critical values $u$ (for $n=4$).
The Puiseux expansions of the critical values $u$ with respect to $\mathsf{T}$ start as follows: for $\lambda>0$, we have
$u$$\sim\begin{cases}(2n+1)n^{-\frac{n}{2n+1}}(q_{1}q_{2}^{n})^{\frac{1}{2n+1}}\mathsf{T}^{\frac{n}{2n+1}\lambda}&\text{($2n+1$ solutions)}\\ (-n)^{n}q_{1}&\text{(one solution)}\end{cases}$
and for $\lambda<0$, we have
(19) $u$ $\sim 2q_{2}^{\frac{1}{2}}\mathsf{T}^{\frac{\lambda}{2}}+(n+1)(q_{1}q_{2}^{\frac{n}{2}})^{\frac{1}{n+1}}\mathsf{T}^{\frac{n}{2n+2}\lambda}\quad\ \text{($2n+2$ solutions)}.$) .
They give two pictures for the critical values (or the eigenvalues of $(c_{1}(X_{n})\star_{\mathbf{q}})$) near the boundaries $\mathcal{C}_{\pi}$, $\mathcal{C}_{\varphi}$. See Figures 5 and 6 for these pictures when $n$ is even/odd.
$\varphi$-picture ($|q_{2}/q_{1}^{2}|$ small)$\pi$-picture ($|q_{2}/q_{1}^{2}|$ large)11112222333344445555666677778888
$\approx$
[…]
Figure 5. Critical values of $f_{q_{1},q_{2}}$ for $n=4$. We let $q_{1},q_{2}$ be positive real numbers and vary $q_{2}/q_{1}^{2}$; the critical values marked by the same number correspond to each other.
$\varphi$-picture ($|q_{2}/q_{1}^{2}|$ small)$\pi$-picture ($|q_{2}/q_{1}^{2}|$ large)11112222333344445555101010109999888877776666
$\approx$
[…]
Figure 6. Two pictures of critical values of $f_{q_{1},q_{2}}$ for $n=5$.
Suppose that $q_{1},q_{2}$ are positive real. The same argument as in Section 3.2 shows that $f_{q_{1},q_{2}}$ has precisely two real critical points, and one of them is the conifold point (whose coordinates are all positive real). Let $T_{-},T_{+}$ denote the critical values of the two real critical points, where $T_{+}=T_{\text{con}}$ corresponds to the conifold point. These values move on the real line and their behaviour (as $q_{2}/q_{1}^{2}$ varies) depends on the parity of $n$; when $n$ is even, they must collide for some value of $q_{2}/q_{1}^{2}$ (see Figures 5, 6). We observe that the other critical values never cross the real line as far as $q_{1},q_{2}$ are positive real, i.e.,
Lemma 7.1.
When $q_{1}$ and $q_{2}$ are positive real, $T_{\pm}$ are the only real critical values of $f_{q_{1},q_{2}}$.
Proof.
By the conformal symmetry (16), it suffices to consider the case where $q_{1}=1$. As discussed in Section 3.2, the equation (17) has precisely two real solutions $t_{\pm}$ and they give rise to the two real critical points. We need to show that non-real solutions of (17) gives rise to a non-real critical value $u$ given by (18). Suppose on the contrary that we have a non-real solution $t$ of (17) (with $\lambda=0$) such that the corresponding $u$ is real. We may assume that $t$ is in the upper-half plane and set $t=re^{i\theta}$ with $r>0$, $\theta\in(0,\pi)$. We have the equalities
$2t^{n+1}+(2n+1)t^{n}=u=2\overline{t}^{n+1}+(2n+1)\overline{t}^{n},$
$t^{2n+2}+nt^{2n+1}=q_{2}=\overline{t}^{2n+2}+n\overline{t}^{2n+1}$
that imply that
(20) $\sin((n+1)\theta)=-\frac{2n+1}{2r}\sin(n\theta),\quad\sin((2n+2)\theta)=-\frac{n}{r}\sin((2n+1)\theta).$
Eliminating $r$, we obtain:
$\sin((n+1)\theta)\sin((2n+1)\theta)=\frac{2n+1}{2n}\sin(n\theta)\sin((2n+2)\theta).$
If $\sin((n+1)\theta)=0$, $\sin(n\theta)=0$ by the equation (20); $\sin((n+1)\theta)=\sin\theta\cos(n\theta)+\cos\theta\sin(n\theta)$ implies that $\cos(n\theta)=0$, which contradicts $\sin(n\theta)=0$. Hence $\sin((n+1)\theta)\neq 0$ and we get
$\sin((2n+1)\theta)=\frac{2n+1}{n}\sin(n\theta)\cos((n+1)\theta).$
We claim that this cannot hold for $\theta\in(0,\pi)$. For general $\theta$, we have
$\sin((2n+1)\theta)-\frac{2n+1}{n}\sin(n\theta)\cos((n+1)\theta)=\frac{(2n+1)\sin\theta-\sin((2n+1)\theta)}{2n}$
by a product-to-sum formula. It suffices to show that $g_{m}(\theta):=\sin(m\theta)/\sin\theta$ has the absolute value less than $m$ for $\theta\in(0,\pi)$ and $m=2,3,4,\dots$. We can easily show this by induction on $m$ using $g_{m+1}(\theta)=\cos(m\theta)+g_{m}(\theta)\cos\theta$. ∎
Remark 7.2.
Recall that $\varphi$ is a blowup at a unique singular point of ${\mathbb{P}}(1^{n+1},n)$. For a blowup $f\colon\widetilde{X}\to X$ of a smooth variety $X$ along a smooth subvariety $Z$ of codimension $r$, the spectrum of $(c_{1}(\widetilde{X})\star_{q})$ near the $f$-boundary has a similar picture: there is a ‘convergent’ group (corresponding to $QH(X)$) surrounded by $(r-1)$ many ‘divergent’ satellite groups (each of which corresponds to $QH(Z)$), see Example 6.5, [Iri20b, Figures 10, 16] and [Iri23b, Figure 5]. However, in the smooth blowup case (unlike $\varphi\colon X_{n}\to\overline{Y}_{n}$ with even $n$), the satellite groups are centered around $C(-1)^{\frac{1}{r-1}}$ for a $C>0$ and do not contain a positive real eigenvalue for a positive real ${\mathbf{q}}$.
Remark 7.3.
We studied a configuration of critical values of the mirror Landau-Ginzburg potential $f_{q_{1},q_{2}}$, or equivalently, eigenvalues of $(c_{1}(X_{n})\star_{\mathbf{q}})$ as ${\mathbf{q}}$ approaches the boundary of $\mathcal{M}_{X_{n}}$. For a general Fano manifold $X$, the spectrum of $(c_{1}(X)\star_{\mathbf{q}})$ up to overall ${\mathbb{R}}_{>0}$-scaling is parametrized by a point in the reduced Kähler moduli space:
$\mathcal{M}_{X}/\text{(real Euler flow)}\cong(H^{2}(X,{\mathbb{R}})/{\mathbb{R}}c_{1}(X))\times H^{2}(X,{\mathbf{i}}{\mathbb{R}})/H^{2}(X,2\pi{\mathbf{i}}{\mathbb{Z}}).$
We can compactify it by adding a sphere $S=S(H^{2}(X,{\mathbb{R}})/{\mathbb{R}}c_{1}(X))$ at infinity on the real part; the compactified space is homeomorphic to $D^{r-1}\times(S^{1})^{r}$ for $r=\dim H^{2}(X)$. It is interesting to understand the limiting behaviour of the spectrum near the boundary: $S$ is decomposed into cells corresponding to extremal rays and the spectrum near each cell is approximated by that of the corresponding exceptional quantum cohomology. In the example at hand, $H^{2}(X_{n},{\mathbb{R}})/{\mathbb{R}}c_{1}(X_{n})\cong{\mathbb{R}}$ has two boundary points corresponding to the extremal contractions $\pi$ and $\varphi$.
7.3. $\widehat{\Gamma}$-integral flat sections and Lefschetz thimbles
Before identifying the principal asymptotic class, we take a brief detour into toric mirror symmetry. We recall the correspondence between $\widehat{\Gamma}$-integral flat sections for quantum connection of $X_{n}$ and Lefschetz thimbles of the Landaug-Ginzburg mirror $f_{q_{1},q_{2}}$, as demonstrated in [Iri09]. Although our attention will be focused on the target spaces $X_{n}$, similar results hold true for weak-Fano toric DM stacks and their mirrors.
To simplify notation (and to emphasize that the results in this section hold for more general toric varieties), we write $X$ for the toric variety $X_{n}={\mathbb{P}}_{{\mathbb{P}}^{n}}({\mathcal{O}}\oplus{\mathcal{O}}(n))$ and $f_{\mathbf{q}}=f_{q_{1},q_{2}}$ for the Laurent polynomial mirror (15) parametrized by ${\mathbf{q}}\in\mathcal{M}_{X}$. Let $N$ be the complex dimension of $X$. Then $f_{\mathbf{q}}$ is a function on (×$({\mathbb{C}}^{\times})^{N}$. We write $x_{1},\dots,x_{N}$ for the variables of $f_{\mathbf{q}}$ (which were denoted by $x_{1},\dots,x_{n},y_{2}$ in (15)). The mirror of the quantum connection is given by the Brieskorn lattice
$\operatorname{Bri}(f_{{\mathbf{q}}}):=H^{N}\left(\Omega_{({\mathbb{C}}^{\times})^{N}}^{\bullet}[z],zd+df_{{\mathbf{q}}}\wedge\right)=\frac{\Omega^{N}_{({\mathbb{C}}^{\times})^{N}}[z]}{(zd+df_{\mathbf{q}}\wedge)\Omega^{N-1}_{({\mathbb{C}}^{\times})^{N}}[z]}$
equipped with the connection $\nabla^{\text{B}}_{z^{2}\partial_{z}}\colon\operatorname{Bri}(f_{\mathbf{q}})\to\operatorname{Bri}(f_{\mathbf{q}})$
$\nabla^{\text{B}}_{z^{2}\partial_{z}}\left[\phi\frac{d{\mathbf{x}}}{{\mathbf{x}}}\right]=\left[\left(z^{2}\frac{\partial\phi}{\partial z}-f_{\mathbf{q}}\phi-\frac{N}{2}z\phi\right)\frac{d{\mathbf{x}}}{{\mathbf{x}}}\right]$
where $\phi\in{\mathbb{C}}[x_{1}^{\pm 1},\dots,x_{N}^{\pm 1},z]$ and $\frac{d{\mathbf{x}}}{{\mathbf{x}}}:=\frac{dx_{1}\cdots dx_{N}}{x_{1}\cdots x_{N}}$. We have a mirror isomorphism of ${\mathbb{C}}[z]$-modules with meromorphic connections:
(21) $\operatorname{Mir}\colon(\operatorname{Bri}(f_{\mathbf{q}}),\nabla^{\text{B}}_{z^{2}\partial_{z}})\cong(H^{*}(X)[z],\nabla^{\text{A}}_{z^{2}\partial_{z}}|_{\mathbf{q}})$
sending $\left[\frac{d{\mathbf{x}}}{{\mathbf{x}}}\right]$ to the identity class $1$, where $\nabla^{\text{A}}|_{\mathbf{q}}$ stands for the quantum connection (13) at ${\mathbf{q}}$. This isomorphism also preserves a connection in the ${\mathbf{q}}$-direction.
The function $f_{\mathbf{q}}\colon({\mathbb{C}}^{\times})^{N}\to{\mathbb{C}}$ defines a locally trivial fibration of $C^{\infty}$ manifolds away from critical values of $f_{\mathbf{q}}$; the symplectic connection associated with the complete Kähler metric $\frac{{\mathbf{i}}}{2}\sum_{i=1}^{N}d\log x_{i}\wedge d\log\overline{x_{i}}$ on (×$({\mathbb{C}}^{\times})^{N}$ defines a well-defined parallel translate between fibres $f_{\mathbf{q}}^{-1}(u)$ along any path on ${\mathbb{C}}$ avoiding critical values. Let $p\in({\mathbb{C}}^{\times})^{N}$ be a non-degenerate critical point of $f_{\mathbf{q}}$ and let $\gamma\colon[0,\infty)\to{\mathbb{C}}$ be an injective path starting from the critical value $\gamma(0)=f_{\mathbf{q}}(p)$ and avoiding any other critical values. We assume that $\gamma(t)$ is linear near $t=0$ and $t=\infty$. When $\gamma^{\prime}(t)\in{\mathbb{R}}_{>0}e^{{\mathbf{i}}\theta}$ near $t=\infty$, we call such a path $\gamma$ an (infinite) vanishing path in the direction $e^{{\mathbf{i}}\theta}$. To the critical point $p$ and the vanishing path $\gamma$, we associate a non-compact cycle $\Gamma\cong{\mathbb{R}}^{N}$
$\Gamma=\Gamma_{p,\gamma}=\bigcup_{t\in[0,\infty)}C_{t}$
where $C_{t}\subset f_{\mathbf{q}}^{-1}(\gamma(t))$ is the vanishing cycle consisting of points that converge to $p$ by parallel translate along $\gamma(t)$ as $t\to+0$ (and $C_{0}=\{p\}$). When $\gamma$ is a vanishing path in the direction $e^{{\mathbf{i}}\theta}$, we call $\Gamma_{p,\gamma}$ an (infinite) Lefschetz thimble in the direction $e^{{\mathbf{i}}\theta}$. When the image of $\gamma$ is the straight ray $f_{\mathbf{q}}(p)+{\mathbb{R}}_{\geq 0}e^{{\mathbf{i}}\theta}$, $\Gamma$ is the stable manifold of $p$ with respect to the downward gradient flow of the function $\operatorname{Re}(e^{-{\mathbf{i}}\theta}f_{\mathbf{q}})$. The infinite Lefschetz thimble $\Gamma$ gives rise to a solution of $\operatorname{Bri}(f_{\mathbf{q}})$
(22) $\operatorname{Bri}(f_{\mathbf{q}})\ni\left[\Omega\right]\longmapsto(-2\pi z)^{-N/2}\int_{\Gamma}e^{f_{\mathbf{q}}({\mathbf{x}})/z}\Omega$
that intertwines $\nabla^{\text{B}}_{z^{2}\partial_{z}}$ with $z^{2}\partial_{z}$. This oscillatory integral converges when $\operatorname{Re}(ze^{-{\mathbf{i}}\theta})<0$ where $e^{{\mathbf{i}}\theta}$ is the direction of $\Gamma$. A Lefschetz thimble in the direction $e^{{\mathbf{i}}\theta}$ gives an element of the relative homology group
$\operatorname{Lef}_{{\mathbf{q}},e^{{\mathbf{i}}\theta}}:=H_{N}(({\mathbb{C}}^{\times})^{N},\{{\mathbf{x}}:\operatorname{Re}(f_{\mathbf{q}}({\mathbf{x}})e^{-{\mathbf{i}}\theta})\geq M\};{\mathbb{Z}})$
where $M\gg 1$ is a sufficiently large number (the right-hand side does not depend on a sufficiently large $M$). These groups $\operatorname{Lef}_{{\mathbf{q}},e^{{\mathbf{i}}\theta}}$ form a local system over $\mathcal{M}_{X}\times S^{1}$. For a generic ${\mathbf{q}}$, $f_{\mathbf{q}}$ has only non-degenerate critical points $p_{1},\dots,p_{s}$. In this case, we set $u_{i}=f_{\mathbf{q}}(p_{i})$ and take mutually non-intersecting vanishing paths $\gamma_{i}\colon[0,\infty)\to{\mathbb{C}}$ starting from $u_{i}$ and extending in the direction of $e^{{\mathbf{i}}\theta}$ (we assume $\gamma_{i}=\gamma_{j}$ when $u_{i}=u_{j}$). Such a system of paths $\gamma_{1},\dots,\gamma_{s}$ is called a distinguished basis of vanishing paths when they are ordered in such a way that
$\operatorname{Im}(\gamma_{1}(t)e^{-{\mathbf{i}}\theta})\geq\operatorname{Im}(\gamma_{2}(t)e^{-{\mathbf{i}}\theta})\geq\cdots\geq\operatorname{Im}(\gamma_{s}(t)e^{-{\mathbf{i}}\theta})\quad\text{for $t\gg 1$.}$
The corresponding Lefschetz thimbles $\Gamma_{p_{1},\gamma_{1}},\dots,\Gamma_{p_{s},\gamma_{s}}$ form a basis of $\operatorname{Lef}_{{\mathbf{q}},e^{{\mathbf{i}}\theta}}$.
Example 7.4.
Suppose that the parameter ${\mathbf{q}}$ is positive real, i.e.
${\mathbf{q}}\in\mathcal{M}_{X,{\mathbb{R}}}:=\operatorname{Hom}(H_{2}(X,{\mathbb{Z}}),{\mathbb{R}}_{>0})\subset\mathcal{M}_{X}.$
The Lefschetz thimble associated with the conifold point ${\mathbf{x}}_{\text{con}}\in({\mathbb{R}}_{>0})^{N}$, a unique point where $f_{\mathbf{q}}|_{({\mathbb{R}}_{>0})^{N}}$ attains a minimum, and the vanishing path $f_{\mathbf{q}}({\mathbf{x}}_{\text{con}})+{\mathbb{R}}_{\geq 0}$ is given by $\Gamma_{\mathbb{R}}=({\mathbb{R}}_{>0})^{N}$. We call it the positive-real Lefschetz thimble. Note that this is well-defined (i.e. does not contain other critical points in the closure) even when $f_{\mathbf{q}}({\mathbf{x}}_{\text{con}})+{\mathbb{R}}_{\geq 0}$ contains other critical values.
The main result of [Iri09] concerns a relationship between oscillatory integrals (22) associated with Lefschetz thimbles $\Gamma$ and $\widehat{\Gamma}$-integral flat sections for the quantum connection. Let $K^{0}(X)$ denote the $K$-group of topological ${\mathbb{C}}$-vector bundles on $X$. The $\widehat{\Gamma}$-integral flat section $\mathfrak{s}_{V}$ associated with $V\in K^{0}(X)$ is defined to be ([Iri09, Definition 2.9])
$\mathfrak{s}_{V}(\tau,z):=(2\pi)^{-N/2}S(\tau,z)z^{-\mu}z^{c_{1}(X)}\left(\widehat{\Gamma}_{X}\operatorname{Ch}(V)\right)$
where $S(\tau,z)z^{-\mu}z^{c_{1}(X)}$ is the fundamental solution in the previous section §6.2 and $\operatorname{Ch}(V)=\sum_{k\geq 0}(2\pi{\mathbf{i}})^{k}\operatorname{ch}_{k}(V)$ is the modified Chern character. The map $V\mapsto\mathfrak{s}_{V}(\tau,z)$ intertwines the Euler pairing with the Poincaré pairing $\langle\cdot,\cdot\rangle^{X}$ as follows [Iri09, Proposition 2.10]:
$\chi(V_{1},V_{2})=\left\langle\mathfrak{s}_{V_{1}}(\tau,e^{-{\mathbf{i}}\pi}z),\mathfrak{s}_{V_{2}}(\tau,z)\right\rangle^{X}$
where $\chi(V_{1},V_{2})$ is given by $\sum_{k=0}^{N}(-1)^{k}\dim\operatorname{Ext}^{k}(V_{1},V_{2})$ if $V_{1},V_{2}$ are holomorphic vector bundles; otherwise, it is the $K$-theoretic push-forward of $V_{1}^{*}\otimes V_{2}$ to a point. The lattice $\operatorname{Lef}_{{\mathbf{q}},1}$ on the mirror side also has a pairing. The perfect intersection pairing #$\#\colon\operatorname{Lef}_{{\mathbf{q}},-1}\times\operatorname{Lef}_{{\mathbf{q}},1}\to{\mathbb{Z}}$ induces a pairing
#$\#(e^{-\pi{\mathbf{i}}}\Gamma_{1}\cdot\Gamma_{2})=\#(\Gamma_{1}\cdot e^{\pi{\mathbf{i}}}\Gamma_{2})$
of $\Gamma_{1},\Gamma_{2}\in\operatorname{Lef}_{{\mathbf{q}},1}$, where $e^{\pm\pi{\mathbf{i}}}\Gamma\in\operatorname{Lef}_{{\mathbf{q}},-1}$ is the parallel translate of $\Gamma\in\operatorname{Lef}_{{\mathbf{q}},1}$ along the semicircle $[0,\pi]\ni\theta\mapsto({\mathbf{q}},e^{\pm{\mathbf{i}}\theta})\in\mathcal{M}_{X}\times S^{1}$.
Theorem 7.5 ([Iri09, §4.3]).
For a positive real parameter ${\mathbf{q}}=e^{\tau}\in\mathcal{M}_{X,{\mathbb{R}}}$ with $\tau\in H^{2}(X,{\mathbb{R}})$, we have an isomorphism of lattices
(23) $\operatorname{Lef}_{{\mathbf{q}},1}\cong K^{0}(X),\quad\Gamma\mapsto V(\Gamma)$
which is locally constant in ${\mathbf{q}}\in\mathcal{M}_{X,{\mathbb{R}}}$ such that for $z>0$,
(24) $(2\pi z)^{-N/2}\int_{\Gamma}e^{-f_{\mathbf{q}}({\mathbf{x}})/z}\Omega_{-z}=\left\langle\iota^{*}(\operatorname{Mir}[\Omega]),\mathfrak{s}_{V(\Gamma)}(\tau,z)\right\rangle^{X}$
for $[\Omega]=[\Omega_{z}]\in\operatorname{Bri}(f_{\mathbf{q}})$, where $\iota^{*}\colon H^{*}(X)[z]\to H^{*}(X)[z]$ is the map sending $f(z)$ to $f(-z)$. Moreover, we have
(25) $\begin{split}V(\Gamma_{\mathbb{R}})&={\mathcal{O}}_{X}\\ \chi(V(\Gamma_{1}),V(\Gamma_{2}))&=(-1)^{N(N-1)/2}\#(e^{-\pi{\mathbf{i}}}\Gamma_{1}\cdot\Gamma_{2})\end{split}$
for $\Gamma_{1},\Gamma_{2}\in\operatorname{Lef}_{{\mathbf{q}},1}$, where $\Gamma_{\mathbb{R}}$ is the positive-real Lefschetz thimble in Example 7.4.
The correspondence (23) between Lefschetz thimbles and $K$-classes of vector bundles is expected to be induced by homological mirror symmetry: an equivalence between the Fukaya-Seidel category of $f_{\mathbf{q}}$ and the derived category of coherent sheaves on $X$. The recent works [Fan20, FaZh] confirmed this expectation for a version of the Fukaya-Seidel category for toric Fano manifolds $X$, thereby showing the Gamma conjecture II for them. Note that a Lefschetz thimble is an exceptional object in the Fukaya-Seidel category; moreover, a distinguished system of vanishing paths gives rise to a full exceptional collection of Lefschetz thimbles.
Remark 7.6.
We collect several references and facts omitted in the above expositions.
(1) The Brieskorn lattice for a tame function on an affine variety has been studied in [Sab99, DuSa]. The Brieskorn lattice of a convenient and non-degenerate Laurent polynomial $f$ (in the sense of Kouchnirenko [Kou76, Definitions 1.5, 1.19]) is free of rank equal to $\dim(\operatorname{Jac}(f))$. The Landau-Ginzburg potential $f_{\mathbf{q}}$ mirror to a toric Fano manifold is non-degenerate for all ${\mathbf{q}}\in\mathcal{M}_{X}$.
(2) The mirror isomorphism (21) follows from Givental’s mirror theorem [Giv98] for weak-Fano toric manifolds (generalized to toric DM stacks in [CCIT]). A detailed comparison of connections (together with pairings) is given in [Iri09, Proposition 4.8], [ReSe, Proposition 4.10], [Iri20b, Proposition 6.16] for weak-Fano [^4] In the weak-Fano case, in general, the mirror isomorphism holds for ${\mathbf{q}}$ in a small neighbourhood of the origin ${\mathbf{q}}=0$ in ¯$\overline{\mathcal{M}}_{X}$; we also need a non-trivial mirror map between the parameter ${\mathbf{q}}$ of the Landau-Ginzburg potentials $f_{\mathbf{q}}$ and the Kähler parameter ${\mathbf{q}}\in\mathcal{M}_{X}$. For toric Fano manifolds, the mirror map is the identity map and the mirror isomorphism holds globally. toric DM stacks/manifolds.
(3) The symplectic connection defines a well-defined $C^{\infty}$ parallel translate between regular fibres $f_{\mathbf{q}}^{-1}(u)$ when $f_{\mathbf{q}}$ is a convenient and non-degenerate Laurent polynomial. This follows from the proof of [Iri20b, Corollary 7.21].
(4) That the map $\Gamma\mapsto V(\Gamma)$ preserves the pairing follows from the fact that the mirror isomorphism preserves the pairing [Iri09, §3.3.1-3.3.2, Proposition 4.8]. The sign factor $(-1)^{N(N-1)/2}$ was missing in [Iri09] and corrected in [Iri11, footnote (16)].
Since $H^{2}(X,{\mathbb{C}})$ is a universal covering of $\mathcal{M}_{X}$, we can uniquely extend the isomorphism (23) to any $\tau\in H^{2}(X,{\mathbb{C}})$
(26) $\operatorname{Lef}_{e^{\tau},1}\cong K^{0}(X),\quad\Gamma\mapsto V(\Gamma)$
so that it is locally constant in $\tau$.
Theorem 7.7.
Suppose that $(c_{1}(X)\star_{\mathbf{q}})$ has a simple rightmost eigenvalue $u$ for some ${\mathbf{q}}=e^{\tau}\in\mathcal{M}_{X}$. Then $u$ equals $f_{\mathbf{q}}(p)$ for a unique non-degenerate critical point $p$ of $f_{\mathbf{q}}$ and $\gamma\colon[0,\infty)\to{\mathbb{C}}$, $\gamma(t)=u+t$ gives a vanishing path. The principal asymptotic class at $\tau$ is given by
$A_{X}(\tau)=\widehat{\Gamma}_{X}\operatorname{Ch}(V(\Gamma_{p,\gamma}))$
where $V(\Gamma_{p,\gamma})$ is given by the isomorphism (26) extended to $\tau\in H^{2}(X,{\mathbb{C}})$. In particular, if $\tau\in H^{2}(X,{\mathbb{R}})$ and $p$ is the conifold point (such that $u=f_{\mathbf{q}}(p)$ is a simple rightmost eigenvalue), we have $A_{X}(\tau)=\widehat{\Gamma}_{X}$.
Proof.
The statement about $p$ and $\gamma$ follows from the isomorphism $QH(X)|_{\mathbf{q}}\cong\operatorname{Jac}(f_{\mathbf{q}})$ sending $c_{1}(X)$ to $[f_{\mathbf{q}}]$. By analytic continuation in $\tau$, we obtain from (24) that
$\left\langle\iota^{*}(\operatorname{Mir}[\Omega]),S(\tau,z)z^{-\mu}z^{c_{1}(X)}\widehat{\Gamma}_{X}\operatorname{Ch}(V(\Gamma_{p,\gamma}))\right\rangle^{X}=z^{-N/2}\int_{\Gamma_{p,\gamma}}e^{-f_{\mathbf{q}}({\mathbf{x}})/z}\Omega_{-z}$
holds for all $\Omega\in\Omega^{N}_{({\mathbb{C}}^{\times})^{N}}[z]$ where ${\mathbf{q}}=e^{\tau}$. The right-hand side has the asymptotic expansion of the form $e^{-u/z}(a_{0}+a_{1}z+\cdots)$ as $z\to+0$ and therefore the flat section
$e^{u/z}S(\tau,z)z^{-\mu}z^{c_{1}(X)}\widehat{\Gamma}_{X}\operatorname{Ch}(V(\Gamma_{p,\gamma}))$
is of moderate growth as $z\to+0$ (see the proof of Theorem 4.1). The conclusion follows from the definition of the principal asymptotic class. ∎
7.4. Identification of the principal asymptotic class via mutation
Finally we determine the principal asymptotic class of $X_{n}$ in certain regions of the Kähler moduli space.
The exponential map defines an isomorphism $H^{2}(X,{\mathbb{R}})\xrightarrow{\cong}\mathcal{M}_{X,{\mathbb{R}}}$ onto the positive real locus. We write ${\mathbf{q}}\mapsto\log{\mathbf{q}}\in H^{2}(X,{\mathbb{R}})$ for the inverse map defined for a positive real ${\mathbf{q}}\in\mathcal{M}_{X,{\mathbb{R}}}$.
Recall from §7.2 that the critical value $T_{+}=T_{\text{con}}$ associated with the conifold point gives a simple rightmost eigenvalue of $(c_{1}(X)\star_{\mathbf{q}})$ when $q_{2}/q_{1}^{2}\gg 1$ i.e. near the $\pi$-boundary. When $n$ is odd, the same holds also when $0<q_{2}/q_{1}^{2}\ll 1$, i.e. near the $\varphi$-boundary. Theorem 7.7 immediately implies:
Theorem 7.8.
Suppose that $q_{1},q_{2}$ are positive real. We have $A_{X_{n}}(\log{\mathbf{q}})=\widehat{\Gamma}_{X_{n}}$ when $q_{2}/q_{1}^{2}$ is sufficiently large and $n\geq 1$ is arbitrary, or when $q_{2}/q_{1}^{2}$ is sufficiently small and $n\geq 1$ is odd.
This proposition is a consequence of the fact that the positive real Lefschetz thimble $\Gamma_{\mathbb{R}}$ corresponds to the structure sheaf (Theorem 7.5), i.e.
(27) $(2\pi z)^{-N/2}\int_{\Gamma_{\mathbb{R}}}e^{-f_{\mathbf{q}}({\mathbf{x}})/z}\Omega_{-z}=\left\langle\iota^{*}(\operatorname{Mir}[\Omega]),\mathfrak{s}_{{\mathcal{O}}}(\log{\mathbf{q}},z)\right\rangle^{X_{n}}$
for ${\mathbf{q}}\in\mathcal{M}_{X_{n},{\mathbb{R}}}$ and $z>0$. Using monodromy transformations, we can determine $V(\Gamma)$ for other Lefschetz thimbles in the region $q_{2}/q_{1}^{2}\gg 1$. Suppose now that $n$ is even and write $n=2m$. We write $H:=D_{n+2}$; $H$ is the pull-back of the ample class ${\mathcal{O}}(n)$ on $\overline{Y}_{n}={\mathbb{P}}(1^{n+1},n)$. It is easy to see that the monodromy transformations for $\widehat{\Gamma}$-integral flat sections along the paths $[0,2\pi]\ni\theta\mapsto(e^{{\mathbf{i}}\theta}q_{1},q_{2})$, $[0,2\pi]\ni\theta\mapsto(q_{1},e^{{\mathbf{i}}\theta}q_{2})$ in $\mathcal{M}_{X_{n}}$ correspond respectively to tensoring ${\mathcal{O}}(-D_{1})$ and ${\mathcal{O}}(-H)$ on the $K$-group:
(28) $\begin{split}\mathfrak{s}_{V}(\log q_{1}+2\pi{\mathbf{i}},\log q_{2},z)&=\mathfrak{s}_{V\otimes{\mathcal{O}}(-D_{1})}(\log q_{1},\log q_{2},z),\\ \mathfrak{s}_{V}(\log q_{1},\log q_{2}+2\pi{\mathbf{i}},z)&=\mathfrak{s}_{V\otimes{\mathcal{O}}(-H)}(\log q_{1},\log q_{2},z).\end{split}$
Let $\Gamma_{k}\in\operatorname{Lef}_{{\mathbf{q}},1}$, $k=1,\dots,m$ be the monodromy transforms of $\Gamma_{\mathbb{R}}\in\operatorname{Lef}_{{\mathbf{q}},1}$ along the path $[0,2\pi]\ni\theta\mapsto(e^{{\mathbf{i}}k\theta}q_{1},q_{2})$; the Puiseux expansion (19) shows that they correspond to the vanishing paths starting from the $k$th node depicted in Figure 7. By applying monodromy transformations to (27) and using (28), we have
$(2\pi z)^{-N/2}\int_{\Gamma_{k}}e^{-f_{\mathbf{q}}(x)/z}\Omega_{-z}=\left\langle\iota^{*}(\operatorname{Mir}[\Omega]),\mathfrak{s}_{{\mathcal{O}}(-kD_{1})}(\log{\mathbf{q}},z)\right\rangle^{X_{n}}$
for all $\Omega\in\Omega^{N}_{({\mathbb{C}}^{\times})^{N}}[z]$; it follows that $V(\Gamma_{k})={\mathcal{O}}(-kD_{1})$ under the isomorphism (23). Similarly, we consider the monodromy transform of $\Gamma_{\mathbb{R}}$ along the path $[0,2\pi]\ni\theta\mapsto(e^{-{\mathbf{i}}m\theta}q_{1},e^{{\mathbf{i}}\theta}q_{2})$; it is a Lefschetz thimble $\widetilde{\Gamma}_{-}$ associated with the ‘greatly bent’ vanishing path starting from $T_{-}$ as in Figure 7. We find that $V(\widetilde{\Gamma}_{-})={\mathcal{O}}(-H+mD_{1})$.
[…]
Figure 7. Vanishing paths and corresponding coherent sheaves (when $q_{2}/q_{1}^{2}$ large and $n=4$): thimbles of solid paths are all given by monodromy transformations from the positive-real Lefschetz thimble $\Gamma_{\mathbb{R}}$.
We want to find the class of the Lefschetz thimble $\Gamma_{-}$ corresponding to the red dashed vanishing path from $T_{-}$ (straight ray in the direction $e^{{\mathbf{i}}\varepsilon}$ with $0<\varepsilon\ll 1$) in Figure 7 using mutation (or the Picard-Lefschetz transformation). Let us first recall the rudiments of mutation for Lefschetz thimbles or (complexes of) coherent sheaves. Figure 8 shows an example of the right mutation $R_{\Delta}\Gamma$ of a Lefschetz thimble $\Gamma$ with respect to another one $\Delta$. By the Picard-Lefschetz monodromy formula (see e.g. [AGV]), we find that the relative homology class of $\Gamma$ equals the sum of $k\Delta$ and $R_{\Delta}\Gamma$, where $k$ is given by the intersection number of vanishing cycles $C_{\Gamma}$, $C_{\Delta}$ associated with $\Gamma,\Delta$:
$k=(-1)^{(N-1)(N-2)/2}\#(C_{\Gamma}\cdot C_{\Delta})=(-1)^{N(N-1)/2}\#(e^{-\pi{\mathbf{i}}}\Gamma\cdot\Delta).$
Therefore we have
$[R_{\Delta}\Gamma]=[\Gamma]-(-1)^{N(N-1)/2}\#(e^{-\pi{\mathbf{i}}}\Gamma\cdot\Delta)[\Delta].$
When we write $e,f,R_{e}f\in K^{0}(X)$ for the $K$-classes corresponding to $\Delta,\Gamma,R_{\Delta}\Gamma$ under the isomorphism (23), we have
$R_{e}f=f-\chi(f,e)e$
by (25). We define the right mutation $R_{e}f$ of $f\in K^{0}(X)$ with respect to $e\in K^{0}(X)$ by this formula. This operation can be lifted to the derived category. The right mutation of an object ${\mathcal{F}}\in D^{b}(X)$ with respect to $\mathcal{E}\in D^{b}(X)$ is defined to be
$R_{\mathcal{E}}{\mathcal{F}}=\operatorname{Cone}({\mathcal{F}}\to\operatorname{Hom}^{\bullet}({\mathcal{F}},\mathcal{E})^{\vee}\otimes\mathcal{E})[-1]$
where $\operatorname{Hom}^{\bullet}({\mathcal{F}},\mathcal{E})^{\vee}\otimes\mathcal{E}=\bigoplus_{k}\operatorname{Hom}^{k}({\mathcal{F}},\mathcal{E})^{\vee}\otimes\mathcal{E}[k]$ and ${\mathcal{F}}\to\operatorname{Hom}^{\bullet}({\mathcal{F}},\mathcal{E})^{\vee}\otimes\mathcal{E}$ is the co-evaluation map. This operation preserves exceptional collections, i.e. if $(\mathcal{E}_{1},\dots,\mathcal{E}_{i},\mathcal{E}_{i+1},\dots,\mathcal{E}_{n})$ is an exceptional collection in $D^{b}(X)$, then so is $(\mathcal{E}_{1},\dots,\mathcal{E}_{i+1},R_{\mathcal{E}_{i+1}}\mathcal{E}_{i},\dots,\mathcal{E}_{n})$. We refer the reader to [BoPo] for the details.
[…]
Figure 8. Right mutation of Lefschetz thimbles.
Lemma 7.9.
Suppose that $n=2m$ is even. In the derived category $D^{b}(X_{n})$ of coherent sheaves, we have
$R_{{\mathcal{O}}(-D_{1})}R_{{\mathcal{O}}(-2D_{1})}\cdots R_{{\mathcal{O}}(-mD_{1})}{\mathcal{O}}(-H+mD_{1})={\mathcal{O}}_{E}(-m)[-1]$
where $E:=D_{n+3}\cong{\mathbb{P}}^{n}$ is the exceptional divisor of $\varphi\colon X_{n}\to\overline{Y}_{n}$. Moreover, this object is orthogonal [^5] Two objects $A,B\in D^{b}(X)$ are said to be orthogonal if $\operatorname{Hom}^{\bullet}(A,B)=\operatorname{Hom}^{\bullet}(B,A)=0$. to ${\mathcal{O}}(-jD_{1})$ when $|j|\leq m-1$.
Proof.
We begin with computing $R_{{\mathcal{O}}(-mD_{1})}{\mathcal{O}}(-H+mD_{1})$. We have
$\operatorname{Hom}^{\bullet}({\mathcal{O}}(-H+mD_{1}),{\mathcal{O}}(-mD_{1}))=H^{\bullet}(X_{n},{\mathcal{O}}(-2mD_{1}+H))=H^{\bullet}(X_{n},{\mathcal{O}}(E))$
by $E=D_{n+3}=D_{n+2}-nD_{1}$. Using the long exact sequence associated with
(29) $0\to{\mathcal{O}}\to{\mathcal{O}}(E)\to{\mathcal{O}}_{E}(E)\cong{\mathcal{O}}_{{\mathbb{P}}^{n}}(-n)\to 0$
and the Kodaira vanishing for ${\mathcal{O}}_{{\mathbb{P}}^{n}}(-n)={\mathcal{O}}_{{\mathbb{P}}^{n}}(1)\otimes K_{{\mathbb{P}}^{n}}$, we get $H^{\bullet}(X_{n},{\mathcal{O}}(E))=H^{\bullet}(X_{n},{\mathcal{O}})={\mathbb{C}}$. Hence
$R_{{\mathcal{O}}(-mD_{1})}{\mathcal{O}}(-H+mD_{1})$ $=\operatorname{Cone}({\mathcal{O}}(-H+mD_{1})\to{\mathcal{O}}(-mD_{1}))[-1]$
$\cong{\mathcal{O}}(-mD_{1})\otimes\operatorname{Cone}({\mathcal{O}}(-E)\to{\mathcal{O}})[-1]$
$\cong{\mathcal{O}}(-mD_{1})\otimes{\mathcal{O}}_{E}[-1]\cong{\mathcal{O}}_{E}(-m)[-1].$
We denote this object by $\mathcal{E}$. We claim that $\mathcal{E}$ is orthogonal to ${\mathcal{O}}(-jD_{1})$ when $|j|\leq m-1$. Applying $\operatorname{Hom}^{\bullet}(-,{\mathcal{O}}(-jD_{1}))$ to the triangle
${\mathcal{O}}(-E-mD_{1})\to{\mathcal{O}}(-mD_{1})\to\mathcal{E}[1],$
we obtain the long exact sequence
(30) $\operatorname{Hom}^{\bullet}(\mathcal{E}[1],{\mathcal{O}}(-jD_{1}))\to H^{\bullet}(X_{n},{\mathcal{O}}((m-j)D_{1}))\to H^{\bullet}(X_{n},{\mathcal{O}}(E+(m-j)D_{1}))\to\cdots.$
Using the long exact sequence associated with (29) tensored by ${\mathcal{O}}((m-j)D_{1})$ and the Kodaira vanishing for ${\mathcal{O}}_{E}(E+(m-j)D_{1})\cong{\mathcal{O}}_{{\mathbb{P}}^{n}}(-(m+j))\cong{\mathcal{O}}_{{\mathbb{P}}^{n}}(m-j+1)\otimes K_{{\mathbb{P}}^{n}}$, we find that the map
$H^{\bullet}(X_{n},{\mathcal{O}}((m-j)D_{1}))\to H^{\bullet}(X_{n},{\mathcal{O}}((m-j)D_{1}+E))$
induced by the natural inclusion ${\mathcal{O}}\hookrightarrow{\mathcal{O}}(E)$ is an isomorphism. Therefore $\operatorname{Hom}^{\bullet}(\mathcal{E}[1],{\mathcal{O}}(-jD_{1}))=0$ by (30). Also we have
$\operatorname{Hom}^{\bullet}({\mathcal{O}}(-jD_{1}),\mathcal{E})$ $\cong H^{\bullet-1}({\mathbb{P}}^{n},{\mathcal{O}}_{{\mathbb{P}}^{n}}(-(m-j)))=0$
again by the Kodaira vanishing. The claim follows. It implies that
$R_{{\mathcal{O}}(-D_{1})}\cdots R_{{\mathcal{O}}(-(m-1)D_{1})}\mathcal{E}=\mathcal{E}$
and the conclusion follows. ∎
Because the mutation of Lefschetz thimbles corresponds to the mutation of $K$-classes, the above lemma shows that $V(\Gamma_{-})=-[{\mathcal{O}}_{E}(-m)]$ when $q_{2}/q_{1}^{2}$ is large. Furthermore, since this class is orthogonal to ${\mathcal{O}}$, the classes of the Lefschetz thimbles associated with the vanishing paths $T_{-}+e^{{\mathbf{i}}\epsilon}{\mathbb{R}}_{\geq 0}$ and $T_{-}+e^{-{\mathbf{i}}\epsilon}{\mathbb{R}}_{\geq 0}$ are the same for a sufficiently small $\epsilon>0$.
Now we vary ${\mathbf{q}}$ within the positive real locus and pass from the $\pi$-picture ($q_{2}/q_{1}^{2}$ large) to the $\varphi$-picture ($q_{2}/q_{1}^{2}$ small). Due to orthogonality, there is no interaction (mutation) between the thimbles $\Gamma_{-}$ and $\Gamma_{\mathbb{R}}$ associated with $T_{-}$ and $T_{+}=T_{\text{con}}$ respectively, even though $T_{\pm}$ collide. Lemma 7.1 guarantees no interactions with other critical points either. Therefore we have $V(\Gamma_{-})=-[{\mathcal{O}}_{E}(-m)]$ for all $q_{1},q_{2}>0$ and conclude the following by Theorem 7.7:
Theorem 7.10.
Let $n\geq 2$ be even. Suppose that $q_{1},q_{2}$ are positive real and that $T_{-}$ is a simple rightmost eigenvalue of $(c_{1}(X_{n})\star_{\mathbf{q}})$; this happens when $q_{2}/q_{1}^{2}$ is sufficiently small or when $n\geq 4$ and $q_{2}/q_{1}^{2}=1$. Then the principal asymptotic class is given by
$A_{X_{n}}(\log{\mathbf{q}})=\widehat{\Gamma}_{X_{n}}\operatorname{Ch}({\mathcal{O}}_{E}(-m)).$
Remark 7.11.
The principal asymptotic class in Theorem 7.10 is a purely imaginary class, confirming the latter half of Corollary 6.11. The limit
$\lim_{t\to\infty}\frac{J_{X}(c_{1}\log t,1)}{\langle J_{X}(c_{1}\log t,1),[\operatorname{pt}]\rangle^{X}}$
appearing in the original Gamma conjecture I [GGI, Corollary 3.6.9] does not exist in this case, as $A_{X_{n}}$ is supported on the exceptional divisor and $\langle A_{X_{n}},[\operatorname{pt}]\rangle^{X_{n}}$ vanishes. Instead, we need to consider the limit in the projective space ${\mathbb{P}}(H^{*}(X))$ as in Proposition 6.10.
Using monodromy in the $\pi$-picture and passing to the $\varphi$-picture, we can determine the $K$-classes corresponding to a distinguished basis of vanishing paths. See Figure 9 for the $n=4$ case.
[…]
Figure 9. A distinguished basis of vanishing paths and the corresponding objects in $D^{b}(X_{n})$, when $q_{2}/q_{1}^{2}$ is small and $n=4$.
Appendix A Conjecture $\mathcal{O}$ for $\mathbb{P}_{\mathbb{P}^{n}}(\mathcal{O}\oplus\mathcal{O}(n-1))$.
In the appendix, we investigate Conjecture $\mathcal{O}$ for $X_{\Delta}=\mathbb{P}_{\mathbb{P}^{n}}(\mathcal{O}\oplus\mathcal{O}(n-1))$, which is again a toric Fano manifold of Picard number two and of Fano index two.
When $n=1$, $X_{\Delta}=\mathbb{P}^{1}\times\mathbb{P}^{1}$. When $n=2$, $X=Bl_{\mathrm{pt}}\mathbb{P}^{3}$ is the blow-up of $\mathbb{P}^{3}$ of a point, and has been studied in [Yang]. Now we assume $n\geq 3$. The arguments are similar to and slightly more involved than that for $X_{n}$. We include the details for two reasons: (a) Property $\mathcal{O}$ for $X_{\Delta}$ of higher Fano index can also be addressed with this method; (b) the eigenvalues of $\hat{c}_{1}$ of maximum modulo could all belong to $\mathbb{C}\setminus\mathbb{R}$.
The mirror superpotential $f$ of $X_{\Delta}$ is given by
(31) $f(\textbf{x})=x_{1}+x_{2}+\cdots+x_{n}+x_{n+1}+\frac{(x_{n+1})^{n-1}}{x_{1}x_{2}\cdots x_{n}}+\frac{1}{x_{n+1}}.$
By solving $\frac{\partial f}{\partial x_{i}}=0$, we have $x:=x_{1}=\cdots=x_{n}$, $y:=x_{n+1}$, $x,y\in\mathbb{C}^{\times}$,
$y^{n-1}=x^{n+1},\quad 1+(n-1)\frac{y^{n-2}}{x^{n}}-\frac{1}{y^{2}}=0.$
Case: $n$ is odd, with $m:=\frac{n-1}{2}\geq 1$
We have $(x^{m+1}+y^{m})(x^{m+1}-y^{m})=0$, resulting in the parameterization $(x,y)=\pm(t^{m},t^{m+1})$ together with the constraint
$h_{1}(t):=t^{2m+2}+(2m)t^{2m+1}=1.$
At every root $t$ of $h_{1}(t)-1$, we have $g_{1}(t)=\tilde{g}_{1}(t)=\check{g}_{1}(t)$, where
$g_{1}(t):=(2m+1)t^{m}+t^{m+1}+t^{m}+\frac{1}{t^{m+1}},\quad\tilde{g}_{1}(t):=2t^{m}+\frac{2}{t^{m+1}},\quad\check{g}_{1}(t):=2t^{m+1}+(4m+2)t^{m}.$
Moreover, all the $2n+2=2(2m+2)$ critical values of $f$ (with multiplicity counted) are precisely given by ±$\pm g_{1}(t)$ at these roots. That is, it suffices to consider the following optimization problem in nonlinear programming.
Problem A.1.
$\textnormal{Maximize}\quad\big{|}g_{1}(t)\big{|}\quad\textnormal{subject to}\quad h_{1}(t)=1,\quad t\in\mathbb{C}^{\times}.$
With similar arguments to $h(t)$ for $X_{n}$, we see that every root of $h_{1}(t)-1$ is of multiplicity one, and $h_{1}(t)-1$ has exactly two simple real roots $a_{1\pm}$ with
$a_{1+}\in(0,1),\qquad a_{1-}\in(-2m-1,-2m).$
Noting the critical points of $f$ are of the form ±$\pm(t^{m},\cdots,t^{m},t^{m+1})$, for $f$ we have
Proposition A.2.
$\mathbf{x}_{\rm con}=(a_{1+}^{m},\cdots,a_{1+}^{m},a_{1+}^{m+1})$, hence $T_{\rm con}=g_{1}(a_{1+})$.
Proposition A.3.
All the roots of $h_{1}(t)-1$ but $a_{1-}$ are in the annulus $N_{[a_{1+},1]}$. Moreover, we have $\tilde{g}_{1}(a_{1+})=\max_{t\in N_{[a_{1+},1]}}|\tilde{g}_{1}(t)|.$
Proof.
By comparing with $(2m)t^{2m+1}$ and using Rouché’s theorem, we conclude that $h_{1}(t)-1$ has $12m+$ 1 zeros in $D_{1}$. By the maximum modulus principle, $\mathring{D}_{a_{1+}}$ contains no zeros of $h_{1}(t)-1$. The first statement follows by noting $a_{1-}\in(-2m-1,-2m)$. The argument for the second statement is the same as that for Lemma 3.7. ∎
The nonlinear programming Problem A.1 can be answered by the following.
Proposition A.4.
Let $n$ be odd with $m=\frac{n-1}{2}\geq 1$. Take $\alpha\in\mathbb{C}\setminus\{a_{1+},a_{1-}\}$ with $h_{1}(\alpha)=1$. We have
  1. (1)
    $|g_{1}(\alpha)|<g_{1}(a_{1+})<4m+4$.
  2. (2)
    $2(2m)^{m}-1<(-1)^{m}g_{1}(a_{1-})<2(2m)^{m}$.
  3. (3)
    If $m\geq 2$, then $|g_{1}(a_{1-})|>g_{1}(a_{1+})$; if $m=1$, $g_{1}(a_{1+})>|g_{1}(a_{1-})|$.
Proof.
Statement (1) follows by the same arguments as in Proposition 3.8.
Viewing $h_{1},\check{g}_{1}$ as real functions in $[-2m-1,-2m]$, we have
$0<(-1)^{m}\check{g}_{1}^{\prime}(x)=(-1)^{m}((2m+2)x+4m^{2}+2m)x^{m-1}<-((2m+2)x+4m^{2}+2m)x^{2m}=-h_{1}^{\prime}(x).$
Hence, $(-1)^{m}g_{1}(a_{1-})=(-1)^{m}\check{g}_{1}(a_{1-})<(-1)^{m}\check{g}_{1}(-2m)=2(2m)^{m}$ and $(-1)^{m}g_{1}(a_{1-})>(-1)^{m}\check{g}_{1}(-2m)-1=2(2m)^{m}-1$, by noting
[…]
Thus statement (2) holds.
For $m\geq 2$, $|g_{1}(a_{1-})|>2(2m)^{m}-1>4m+4>g_{1}(a_{1-})$. For $m=1$, we have $a_{1+}>{7\over 10}$ and
$g_{1}(a_{1+})=\check{g}_{1}(a_{1+})>\check{g}_{1}({7\over 10})=2\cdot({7\over 10})^{2}+6\cdot{7\over 10}>2\cdot 2^{1}>|g_{1}(a_{1-})|>2\cdot 2^{1}-1=3.$
Therefore statement (3) holds. ∎
Case: $n$ is even
$n-1$ and $n+1$ are coprime, so $k_{1}(n-1)+k_{2}(n+1)=1$ for some integers $k_{1},k_{2}$. We can take the parameterization $(x,y)=(t^{n-1},t^{n+1})$ (by setting $t=y^{k_{2}}x^{k_{1}}$), resulting in the constraint
$h_{0}(t):=t^{2n+2}+(n-1)t^{2n}=1.$
Every root $\alpha$ of $h_{0}(t)-1$ is of multiplicity one, and $g_{0}(\alpha)=\tilde{g}_{0}(\alpha)=\check{g}_{0}(\alpha)$, where
$g_{0}(t):=nt^{n-1}+t^{n+1}+t^{n-1}+\frac{1}{t^{n+1}},\quad\tilde{g}_{0}(t):=2t^{n-1}+\frac{2}{t^{n+1}},\quad\check{g}_{0}(t):=2t^{n+1}+(2n)t^{n-1}.$
Moreover, all the $(2n+2)$ critical values of $f$, with multiplicity counted, are given by $g_{0}(t)$ at these roots. It suffices to consider the following optimization problem.
Problem A.5.
$\textnormal{Maximize}\quad\big{|}g_{0}(t)\big{|}\quad\textnormal{subject to}\quad h_{0}(t)=1,\quad t\in\mathbb{C}^{\times}.$
With similar arguments to $h_{1}(t)-1$, we see that $u^{n+1}+(n-1)u-1$ consists of two real roots, which are both of multiplicity one and belong to $(0,1)$ and $(-n+1,-n+{3\over 2})$ respectively. Hence, $h_{0}(t)-1$ has four roots ±$\pm a_{0+}$ and ±$\pm a_{\rm I}$, with
$a_{0+}\in(0,1),\qquad a_{\rm I}\in{\mathbf{i}}\mathbb{R}\quad\mbox{and}\quad\mathrm{Im}(a_{\rm I})\in(\sqrt{n-\frac{3}{2}},\sqrt{n-1}).$
Moreover, ±$\pm a_{0+}$ are the only real roots of $h_{0}(t)-1$. Noting the critical points of $f$ are of the form $(t^{n-1},\cdots,t^{n-1},t^{n+1})$, we have
Proposition A.6.
$\mathbf{x}_{\rm con}=(a_{0+}^{n-1},\cdots,a_{0+}^{n-1},a_{0+}^{n+1})$, hence $T_{\rm con}=g_{0}(a_{0+})$.
The following is parallel to Proposition A.3, obtained the same arguments.
Proposition A.7.
All the roots of $h_{0}(t)-1$ but ±$\pm a_{\rm I}$ are in the annulus $N_{[a_{0+},1]}$. Moreover, we have $\tilde{g}_{0}(a_{0+})=\max_{t\in N_{[a_{0+},1]}}|\tilde{g}_{0}(t)|.$
Proposition A.8.
Let $n\geq 4$ be even. Take any $\alpha\in\mathbb{C}\setminus\{\pm a_{0+},\pm a_{\rm I}\}$ with $h_{0}(\alpha)=1$. We have
  1. (1)
    $|g_{0}(\alpha)|<g_{0}(a_{0+})=-g_{0}(-a_{0+})<2n+2$.
  2. (2)
    $12(n-1)^{\frac{n-1}{2}}<{\mathbf{i}}^{n+1}g(a_{\rm I})={\mathbf{i}}^{n-1}g(-a_{\rm I})<2(n-1)^{\frac{n-1}{2}}+$ 1.
Proof.
It is evident that $g_{0}(-t)=-g_{0}(t)$, thus statement (1) follows by the same arguments as in Proposition A.4 (1).
Consider $h_{0}(t),\check{g}_{0}(t)$ as functions on the imaginary axis ${\mathbf{i}}\mathbb{R}$ with $\mathrm{Im}(t)\in[\sqrt{n-\frac{3}{2}},\sqrt{n-1}]$. Denote $v:=\mathrm{Im}(t)$. Taking derivative with respect to $v$, we have
00 0 $>{\mathbf{i}}^{n+1}\check{g}_{0}^{\prime}(v{\mathbf{i}})={\mathbf{i}}^{2n}(2n(n-1)-(2n+2)v^{2})v^{n-2}$
$>{\mathbf{i}}^{2n}(2n(n-1)-(2n+2)v^{2})v^{2n-1}=h_{0}^{\prime}(v{\mathbf{i}}).$
Hence, ${\mathbf{i}}^{n+1}g_{0}(a_{\rm I})={\mathbf{i}}^{n+1}\check{g}_{0}(a_{\rm I})>{\mathbf{i}}^{n+1}\check{g}_{0}(\sqrt{-n+1})=2(n-1)^{\frac{n-1}{2}}$ and ${\mathbf{i}}^{n+1}g_{0}(a_{\rm I})<{\mathbf{i}}^{n+1}\check{g}_{0}(\sqrt{-n+1})+1=2(n-1)^{\frac{n-1}{2}}+1$, by noting
${\mathbf{i}}^{n+1}(\check{g}_{0}(a_{\rm I})-\check{g}_{0}(\sqrt{-n+1}))$
[…]
Thus statement (2) holds. ∎
Theorem A.9.
Conjecture $\mathcal{O}$ holds for $\mathbb{P}_{\mathbb{P}^{n}}(\mathcal{O}\oplus\mathcal{O}(n-1))$ if and only if either $n$ is odd or $n=2$.
Proof.
If $n=1$, then Conjecture $\mathcal{O}$ holds by [HKLY]. If $n=2$, then $\mathbb{P}_{\mathbb{P}^{n}}(\mathcal{O}\oplus\mathcal{O}(1))=Bl_{\rm pt}{\mathbb{P}^{3}}$, Conjecture $\mathcal{O}$ holds by [Yang]. Assume $n\geq 3$ now. By Proposition 2.7 and Propositions A.4, A.8, Property $\mathcal{O}$ (1) holds if and only if $n$ is odd. So does Property $\mathcal{O}$ (2) by noting that $\mathbb{P}_{\mathbb{P}^{n}}(\mathcal{O}\oplus\mathcal{O}(n-1))$ is Fano index two. ∎
Remark A.10.
Our analysis for the superpotential works for more examples. Toric Fano manifolds of Picard number two are classified in [Klei], and are the following projectivization of splitting bundles over projective spaces, $\mathbb{P}_{\mathbb{P}^{s}}\big{(}\mathcal{O}\oplus\bigoplus_{i=1}^{r}\mathcal{O}(a_{i})\big{)},$ where $r,s\geq 1,0\leq a_{1}\leq\cdots\leq a_{r}$, and $\sum_{i=1}^{r}a_{i}\leq s$. Our method can put the analysis for the special case $Bl_{\mathbb{P}^{r-1}}\mathbb{P}^{n}=\mathbb{P}_{\mathbb{P}^{n-r}}(\mathcal{O}(1)\oplus\bigoplus_{i=1}^{r}\mathcal{O})$ in [Yang] uniformly into the framework of nonlinear programming, and works for instance for the toric Fano 4-fold with ID 71 in [Obro2], which of Picard number five.